### v3

# Commitment and Randomization in Communication

Abstract
When does Sender, in a Sender-Receiver game, strictly value commitment? In a setting with finitely many actions and states, we establish that, generically, commitment has no value if and only if a partitional experiment is optimal. Moreover, if Sender’s preferred cheap-talk equilibrium necessarily involves randomization, then Sender values commitment. Our results imply that if a school values commitment to a grading policy, then the school necessarily prefers to grade unfairly. We also ask: for what share of preferences does commitment have no value? For any state space, if there are $\left|A\right|$ actions, the share is at least $\frac{1}{\left|A\right|^{\left|A\right|}}$. As the number of states grows large, the share converges precisely to $\frac{1}{\left|A\right|^{\left|A\right|}}$.
Keywords: Bayesian persuasion; cheap talk
JEL Codes: D80, D83
1 Introduction
Commitment is often valuable. In the context of communication, this fact is brought out by the contrast of Sender’s payoff in Bayesian persuasion versus cheap talk. For any prior, and any profile of Sender and Receiver’s preferences, Sender’s payoff is always weakly higher under Bayesian persuasion than in any cheap-talk equilibrium. [^1] In fact, Bayesian persuasion provides the upper bound on Sender’s equilibrium payoff under any communication protocol, such as disclosure or signaling; see Proposition 1 in the Online Appendix of kamenicagentzkow2011. In this paper, we ask: when does commitment make Sender strictly better off?
Answering this question would contribute to our understanding of circumstances that incentivize building strong institutions that are immune to influence (north1993institutions; lipnowski2022persuasion) or building a reputation for a degree of honesty (best2024persuasion; mathevet2024reputation).
We focus exclusively on environments with finitely many states and actions. We show that, generically, Sender with commitment values that commitment if and only if he values randomization (Theorem 1). [^2] Most of the results in this paper hold generically, i.e., for all but a knife-edge set of preferences. To ease exposition, we henceforth sweep the genericity modifier under the rug in the Introduction. In other words, the Bayesian persuasion payoff is achievable in a cheap-talk equilibrium if and only if a partitional experiment is a solution to the Bayesian persuasion problem. Moreover, if Sender’s preferred equilibrium in a cheap-talk game necessarily involves randomization, then Sender values commitment (Theorem 2).
For an application of these results, consider a school that assigns grades to students, each of whom is characterized by a vector of attributes. Some of the attributes are relevant, in the sense that an employer values those attributes or the school’s value of placing a student depends on them. Other attributes are irrelevant. The school assigns a grade to each student based on her attributes. The school’s grading policy is fair if it assigns the same grade to students with identical relevant attributes. Theorem 1 tells us that if the school values committing to a grading policy of any form (such as mandating a maximum GPA), then the school prefers to grade unfairly. Conversely, if a fair grading scheme is optimal, there is no need for commitment: discretionary “cheap-talk” grades are as effective as those disciplined by a publicly declared grading policy.
We also derive results about the share of preferences such that Sender finds commitment (or, equivalently, randomization) valuable. We uncover a potentially surprising connection between the share of preferences where commitment has value and the cardinalities of the action set and the state space (Theorem 3).
Say that preferences are felicitous if Sender can obtain his ideal payoff under cheap talk. If preferences are felicitous, commitment obviously has no value. We show that, for any number of states, the share of felicitous preferences is at least $\frac{1}{\left|A\right|^{\left|A\right|}}$, where $\left|A\right|$ denotes the cardinality of the action set. Hence, in any state space, the share of preferences such that commitment has no value is at least $\frac{1}{\left|A\right|^{\left|A\right|}}$.
Moreover, as the number of states grows large: (i) the share of felicitous preferences converges to $\frac{1}{\left|A\right|^{\left|A\right|}}$, and (ii) commitment has no value if and only if preferences are felicitous. Thus, when the state space is large, the share of preferences such that commitment has no value is approximately $\frac{1}{\left|A\right|^{\left|A\right|}}$. For example, if the action set is binary and there are many states, commitment is valuable for roughly 75% of possible preferences.
Illustrative example
The workhorse example in the Bayesian-persuasion literature is a prosecutor (Sender) trying to convince a judge (Receiver) to convict a defendant who is guilty or innocent. The judge’s preferences are such that she prefers to convict if the probability of guilt is weakly higher than the probability of innocence. The prosecutor has state-independent preferences and always prefers conviction. The prior probability of guilt is $0.3$.
If the environment is cheap talk, the unique equilibrium outcome is that the judge ignores the prosecutor and always acquits the defendant. If the prosecutor can commit to an experiment about the state, however, he will conduct a stochastic experiment that indicates guilt whenever the defendant is guilty and indicates guilt with probability $\frac{3}{7}$ when the defendant is innocent (kamenicagentzkow2011). This experiment induces the judge to convict the defendant with 60% probability. The prosecutor is thus strictly better off than under cheap talk.
Our Theorem 1 tells us that the two facts, (i) the prosecutor’s optimal experiment involves randomization and (ii) the prosecutor does better under commitment, imply each other. [^3] Theorem 1 only states that (i) and (ii) imply one another for a generic set of preferences. To apply the theorem here, we note that the preferences in the prosecutor-judge example belong to the generic set used in the proof of the Theorem. Moreover, in Online Appendix B.2 we show that Theorem 1 holds when Sender has state-independent preferences. Of course, the prosecutor-judge example was designed to be extremely simple, so in this particular example one can easily determine the optimal experiment and the value of commitment without our result. In more complicated environments, however, Theorem 1 can simplify the determination of whether commitment is valuable. Except in certain cases, such as uniform-quadratic (crawford1982strategic) or transparent preferences (lipnowski2020cheap), cheap-talk games can be difficult to solve. Theorem 1 can then be used to determine whether commitment is valuable without solving for cheap talk equilibria, simply by computing the Bayesian-persuasion optima and checking whether they include a partitional experiment. [^4] Recent research provides a large toolbox for solving Bayesian-persuasion problems, including concavification (kamenicagentzkow2011), price-theoretic approaches (kolotilin2018optimal; dworczak2019simple), duality (dworczak2024persuasion), and optimal-transport theory (kolotilin2023persuasion). bergemann2016bayes show that persuasion problem can be formulated as a linear program; it is well known that linear programs can be computed in polynomial time. In contrast, babichenko2023algorithmic establish that it is NP-hard to approximate Sender’s maximum payoff in cheap-talk, or even to determine if that payoff is strictly greater than in a babbling equilibrium. For a survey of computational approaches to Bayesian persuasion, see dughmi2017algorithmic.
The prosecutor-judge example also illustrates the distinction between the if-and-only-if result in Theorem 1 and the unidirectional Theorem 2. Recall that Theorem 2 does not claim that the value of commitment is positive only if randomization is valuable in cheap talk. The prosecutor-judge example provides a counterexample to such a claim. In the cheap-talk game, the prosecutor has no value for randomization: with or without it, he never obtains any convictions. Yet, the prosecutor obviously values commitment.
Finally, the prosecutor-judge example also helps illustrate what Theorem 1 does not say. Prohibiting randomization would not mean commitment is not valuable. Suppose that the prosecutor is endowed with commitment, but is legally obliged to use only partitional experiments. In that case, the prosecutor would provide a fully informative experiment, obtaining a conviction with 30% probability. That is still better than his cheap-talk payoff of no convictions.
Related literature
Our paper connects the literatures on cheap talk (crawford1982strategic) and Bayesian persuasion (kamenicagentzkow2011). min2021bayesian and lipnowski2022persuasion examine environments with limited commitment that are a mixture of cheap talk and Bayesian persuasion. In contrast, we focus on the question of when cheap talk and Bayesian persuasion yield the same payoff to Sender.555perez2014interim and koessler2023informed examine the circumstances under which Sender attains his Bayesian persuasion payoff even if he learns the state prior to selecting the experiment.
glazer2006study and sher2011credibility consider disclosure games and derive conditions on preferences that imply that Receiver values neither commitment nor randomization. arieli2025communicationleadefficiency establish connections between randomization and efficiency under both cheap talk and Bayesian persuasion.
Several papers examine value of commitment under the assumption that Sender has state-independent preferences. When the action space is finite, as in our framework, lipnowski2020cheap show that (for almost every prior) Sender either: (i) obtains his ideal payoff in cheap talk, or (ii) values commitment; best2024persuasion show that (for almost every prior) Sender either: (i) obtains his ideal payoff under the prior, or (ii) values randomization. titovazhang2024persuasion establish a connection between randomization and the attainability of the Bayesian persuasion payoff under verifiable messages. corrao2023bounds examine Sender’s payoff under cheap talk, mediation, and Bayesian persuasion. They establish that Sender does not value commitment if his payoffs are the same under mediation and Bayesian persuasion.
In the context of mechanism design, value of commitment and value of randomization have been studied separately. Mechanism design with limited commitment has been studied by akbarpourLi and doval2022mechanism, among others. Value of randomization in mechanism design has been widely recognized in single-agent multi-product monopolist settings (e.g., manelli2006bundling). In contrast, with two or more agents, chen2019equivalence establish that if agents’ types are atomless and independently distributed, randomization is never valuable.
2 Set-up and definitions
Preference and beliefs
Receiver (she) has a utility function $u_{R}\left(a,\omega\right)$ that depends on her action $a\in A$ and the state of the world $\omega\in\Omega$. Both $A$ and $\Omega$ are finite; our analysis relies heavily on this assumption. [^6] At the risk of being excessively philosophical, we consider environments with finite $A$ and $\Omega$ to be more realistic; the use of infinite sets often provides tractability but rarely improves realism. We discuss the role of the finiteness assumption in Online Appendix B.1. For any finite set $X$, we denote its cardinality by $\left|X\right|$. Sender (he) has a utility function $u_{S}\left(a,\omega\right)$ that depends on Receiver’s action and the state. The players share an interior common prior $\mu_{0}$ on $\Omega$. We say action $a^{*}$ is $i$’s ideal action in $\omega$ if $a^{*}\in\operatorname*{arg\,max}_{a\in A}u_{i}(a,\omega)$.
Environments, shares, and genericity
We refer to the pair $\left(u_{S},u_{R}\right)$ as the (preference) environment.
Since $u_{S}$ and $u_{R}$ have a finite domain, they are bounded. We further restrict our attention to environments where $u_{S}$ and $u_{R}$ take values in some fixed interval, which, without loss of generality, we set to $\left[0,1\right]$. Under these assumptions, the set of all environments is $\left[0,1\right]^{2\left|A\right|\left|\Omega\right|}$. When we say that a claim holds for a $\gamma$ share of environments, we simply mean that the set of environments where the claim holds has Lebesgue measure $\gamma$ on $\mathbb{R}{}^{2\left|A\right|\left|\Omega\right|}$.
We say a set of environments is generic if it has Lebesgue measure one on $\mathbb{R}{}^{2\left|A\right|\left|\Omega\right|}$. [^7] Our results also hold if we use a topological rather than a measure-theoretic notion of genericity. See Footnotes 23 and 25 in the Appendix. When we say that a claim holds generically, we mean that it holds for a generic set of environments.888lipnowski2020equivalence, who focuses on finite action and state spaces as we do, establishes that commitment has no value when Sender’s value function over Receiver’s beliefs is continuous. Such continuity, however, holds for a zero share of environments. In contrast, we focus on results that hold generically. When we say that, given $A$, a claim holds generically as $\left|\Omega\right|\rightarrow\infty$, we mean that the share of environments where the claim does not hold converges to zero as $\left|\Omega\right|\rightarrow\infty$.
Cheap talk, Bayesian persuasion, and value of commitment
Let $M$ be a finite message space with $|M|>\max\{|\Omega|,|A|\}$. [^9] Our results concern Sender’s payoffs under cheap talk, Bayesian persuasion, and restriction to partitional strategies in those models. To derive Sender’s maximal payoff, it is without loss of generality to set $|M|\geq|\Omega|$ for cheap talk (Matthews90), $|M|\geq\min\{|\Omega|,|A|\}$ for Bayesian persuasion (kamenicagentzkow2011), and $|M|\geq|\Omega|$ for partitional strategies (trivially). Therefore, assuming $|M|\geq|\Omega|$ would suffice for our results. However, further assuming $|M|\geq|A|+1$ simplifies the proofs of Lemmas 4 and 8. Sender chooses a messaging strategy $\sigma:\Omega\rightarrow\Delta M$. Receiver chooses an action strategy $\rho:M\rightarrow\Delta A$.
A profile of strategies $(\sigma,\rho)$ induces expected payoffs
$U_{i}(\sigma,\rho)=\sum_{\omega,m,a}\,\mu_{0}(\omega)\,\sigma(m|\omega)\,\rho(a|m)\,u_{i}(a,\omega)\quad\text{for }i=S,R.$
A profile $(\sigma^{*},\rho^{*})$ is S-BR if $\sigma^{*}\in\operatorname*{arg\,max}_{\sigma}U_{S}(\sigma,\rho^{*})$. A profile $(\sigma^{*},\rho^{*})$ is R-BR if $\rho^{*}\in\operatorname*{arg\,max}_{\rho}U_{R}(\sigma^{*},\rho)$.
Sender’s ideal payoff is the maximum $U_{S}$ induced by any profile.
A cheap-talk equilibrium is a profile that satisfies S-BR and R-BR. [^10] This definition may seem unconventional since it uses Nash equilibrium, rather than perfect Bayesian equilibrium, as the solution concept. In cheap-talk games, however, the set of equilibrium outcomes (joint distributions of states, messages, and actions) is exactly the same whether we apply Nash or perfect Bayesian as the equilibrium concept. The formulation in terms of Nash equilibria streamlines the proofs. We define (Sender’s) cheap-talk payoff as the maximum $U_{S}$ induced by a cheap-talk equilibrium. [^11] Throughout, we examine the value of commitment to Sender; hence the focus on Sender’s payoff. The set of equilibrium payoffs is compact so a maximum exists. We are interested in whether Sender can attain his commitment payoff in some equilibrium, so we focus on Sender-preferred equilibria. Unless no information is the commitment optimum, it cannot be that every cheap-talk equilibrium yields the commitment payoff since every cheap-talk game admits a babbling equilibrium. (And even if no information is the commitment optimum, there might be cheap-talk equilibria that yield a lower payoff to Sender than babbling.)
A persuasion profile is a profile that satisfies R-BR. The (Bayesian) persuasion payoff is the maximum $U_{S}$ induced by a persuasion profile.121212lipnowski2024perfect establish that, with finite $A$ and $\Omega$, this is generically the only payoff that Sender could attain in an equilibrium of a Bayesian persuasion game. We refer to a persuasion profile that yields the persuasion payoff as optimal.
We say that commitment is valuable if the persuasion payoff is strictly higher than the cheap-talk payoff. Otherwise, we say commitment has no value.
Partitional strategies and value of randomization
A messaging strategy $\sigma$ is partitional if for every $\omega$, there is a message $m$ such that $\sigma\left(m|\omega\right)=1$. A profile $\left(\sigma,\rho\right)$ is a partitional profile if $\sigma$ is partitional. [^13] Our focus is on the connection between Sender’s value of commitment and Sender’s randomization. Consequently, the definition of a partitional profile only concerns Sender’s strategy. That said, along the way we will establish a result about Receiver playing pure strategies (see Lemma 4). The partitional persuasion payoff is the maximum $U_{S}$ induced by a partitional persuasion profile. The partitional cheap-talk payoff is the maximum $U_{S}$ induced by a partitional cheap-talk equilibrium. [^14] A partitional cheap-talk equilibrium always exists because the babbling equilibrium outcome can be supported by Sender always sending the same message.
We say that committed Sender values randomization if the persuasion payoff is strictly higher than the partitional persuasion payoff. We say that cheap-talk Sender values randomization if the cheap-talk payoff is strictly higher than the partitional cheap-talk payoff.
3 Value of commitment and randomization under commitment
In this section, we consider a Sender with commitment power, who can choose his messaging strategy prior to being informed of the state. We ask whether this commitment power makes Sender strictly better off. We link the value of commitment to Sender’s behavior under commitment, in particular to whether Sender has a strict preference for randomization.
Theorem 1.
Generically, commitment is valuable if and only if committed Sender values randomization.
Here we provide an intuition about the only-if direction of the theorem. We postpone the discussion of the converse until the next section, as the intuition for it is related to that for Theorem 2. Formal proofs are in the Appendix. [^15] Theorem 1 can be extended to establish a threefold equivalence. Generically, the following imply each other: (i) commitment is valuable, (ii) committed Sender values randomization, and (iii) any optimal persuasion profile induces a belief under which Receiver has multiple optimal actions (see Theorem 1 in the Appendix).
For any $E\subseteq\Omega$, let $\mu_{E}$ denote the posterior belief induced by learning that $\omega$ is in $E$. For a generic set of environments, Receiver’s optimal action given any such $\mu_{E}$ is unique and remains optimal in a neighborhood of beliefs around $\mu_{E}$.
Now, suppose that there is a partitional optimal persuasion profile $\left(\sigma,\rho\right)$. Let $M_{\sigma}$ be the set of messages that are sent under $\sigma$. Because $\sigma$ is partitional, each $m\in M_{\sigma}$ is associated with a subset of the state space, namely $\Omega_{m}\equiv\left\{\omega|\sigma\left(m|\omega\right)=1\right\}$. For each $m\in M_{\sigma}$, let $\mu_{m}$ be the belief induced by $m$, and let $a_{m}$ be Receiver’s (uniquely) optimal action given $\mu_{m}$. As noted above, $a_{m}$ remains optimal in a neighborhood of beliefs around $\mu_{m}$.
Key to the proof is to note that every action $a_{m}$ taken in equilibrium must be Sender’s preferred action, among the actions taken in equilibrium, in all states where action $a_{m}$ is taken. In other words, let $A^{*}=\left\{a_{m}|m\in M_{\sigma}\right\}$; for every $m\in M_{\sigma}$, for every $\omega\in\Omega_{m}$, we have $u_{S}\left(a_{m},\omega\right)\geq u_{S}\left(a_{m^{\prime}},\omega\right)$ for all $a_{m^{\prime}}\in A^{*}$. Why does this hold? If it were not the case, Sender could attain a higher payoff with an alternative strategy: if $u_{S}\left(a_{m},\omega\right)<u_{S}\left(a_{m^{\prime}},\omega\right)$ for some $a_{m^{\prime}}\in A^{*}$, $\omega\in\Omega_{m}$, sender could send $m^{\prime}$ in $\omega$ with a small probability and still keep $a_{m}$ optimal given $m$.
Finally, the fact that for every $m\in M_{\sigma}$, $u_{S}\left(a_{m},\omega\right)\geq u_{S}\left(a_{m^{\prime}},\omega\right)$ for all $a_{m^{\prime}}\in A^{*}$ and all $\omega\in\Omega_{m}$ implies that $\left(\sigma,\rho\right)$ is a cheap-talk equilibrium. [^16] Deviating to an on-path message $\hat{m}\in M_{\sigma}$ cannot be profitable by the inequality $u_{S}\left(a_{m},\omega\right)\geq u_{S}\left(a_{\hat{m}},\omega\right)$ for $\hat{m}\in M_{\sigma}$; for any off-path message $\hat{m}\notin M_{\sigma}$, we can just set $\text{$\rho$$\left(\cdot|\hat{m}\right)$}=\rho\left(\cdot|m^{*}\right)$ for some $m^{*}\in M_{\sigma},$ thus ensuring that such a deviation is also not profitable. Hence, commitment is not valuable.
Theorem 1 only tells us that, generically, commitment has zero value if and only if randomization has zero value. A natural question is whether, generically, small value of commitment implies or is implied by small value of randomization. The answer is no. We construct a positive measure of environments where the value of randomization is arbitrarily small but the value of commitment is not (Online Appendix B.4.1), and a positive measure of environments where the value of commitment is arbitrarily small but the value of randomization is not (Online Appendix B.4.2).
4 Value of commitment and randomization in cheap talk
In this section, we consider a Sender without commitment power who engages in a cheap-talk game. We ask whether he would be strictly better off if he had commitment power. We link the value of such commitment to Sender’s behavior in Sender-preferred cheap-talk equilibria, in particular to whether Sender necessarily randomizes in such equilibria.
Theorem 2.
Generically, commitment is valuable if cheap-talk Sender values randomization.
Theorem 2 and the if-direction of Theorem 1 both derive from the following result. Generically, if a cheap-talk equilibrium yields the persuasion payoff, then there is a partitional $\sigma$ and a (pure-strategy) $\rho$ such that $(\sigma,\rho)$ is a cheap-talk equilibrium and yields the persuasion payoff. We build this result (Proposition 1 in Appendix A.3) in two steps.
The first step (Lemma 4) shows that, generically, if $\left(\sigma,\rho\right)$ is R-BR and yields the persuasion payoff, then $\rho$ must be pure on-path. Consider toward contradiction that there is a message $m$ sent with positive probability under $\sigma$, and there are two distinct actions, say $a$ and $a^{\prime}$, in the support of $\rho\left(\cdot|m\right)$. It must be that both Sender and Receiver are indifferent between $a$ and $a^{\prime}$ under $\mu_{m}$, the belief induced by $m$: Receiver has to be indifferent because $\left(\sigma,\rho\right)$ is R-BR; Sender has to be indifferent because $\left(\sigma,\rho\right)$ yields the persuasion payoff, which maximizes $U_{S}$ over all persuasion profiles. [^17] If Sender strictly prefers one action over the other, say $a$ over $a^{\prime}$, at $\mu_{m}$, then Sender would obtain a higher payoff if Receiver always takes $a$ following $m$ (which would remain R-BR given Receiver’s indifference). The result then follows from establishing that such a coincidence of indifferences generically cannot arise when Sender is optimizing. For some intuition for why this is the case, consider Figure 1 which illustrates this result when there are three states. Suppose $a_{1}$ and $a_{2}$ are in the support of $\rho\left(\cdot|m\right)$. Region $R_{i}$ denotes beliefs where Receiver prefers $a_{i}$. Region $S_{i}$ denotes beliefs where Sender prefers $a_{i}$. Generically, the border between $R_{1}$ and $R_{2}$ is distinct from the border between $S_{1}$ and $S_{2}$ and thus the two borders have at most one intersection, $\mu_{m}$. Moreover, generically $\mu_{m}$ (if it exists) is an interior belief. But now, Sender could deviate to an alternate strategy that induces beliefs $\mu_{1}$ and $\mu_{2}$ instead of $\mu_{m}$, with Receiver still indifferent between $a_{1}$ and $a_{2}$ at both $\mu_{1}$ and $\mu_{2}$. Suppose that Receiver takes action $a_{i}$ following belief $\mu_{i}$. This strategy is still R-BR for Receiver and gives Sender a strictly higher payoff. Thus, we have reached a contradiction. With more than three states and more than two actions, the proof that the coincidence of indifferences generically cannot arise is conceptually similar but notationally more involved. It is presented in the Appendix as Lemma 2.
Figure 1: Indifference incompatible with optimality
The second step (Lemma 5) shows that, generically, if $\left(\sigma,\rho\right)$ is a cheap-talk equilibrium that yields the persuasion payoff, and $\rho$ is a pure strategy on-path, then there is a partitional cheap-talk equilibrium that yields the persuasion payoff. This is easy to see. Generically, for any $\omega$ and any $a\neq a^{\prime}$, we have $u_{S}\left(a,\omega\right)\neq u_{S}\left(a^{\prime},\omega\right)$. Now, consider some cheap-talk equilibrium $\left(\sigma,\rho\right)$, with $\rho$ pure on-path, that yields the persuasion payoff. If $\sigma$ is partitional, our result is immediate. Suppose to the contrary that in some $\omega$, both $m$ and $m^{\prime}$ are sent with positive probability. Then, $m$ and $m^{\prime}$ must induce the same action: if $m$ induces some $a$ and $m^{\prime}$ induces a distinct $a^{\prime}$, the fact that $u_{S}\left(a,\omega\right)\neq u_{S}\left(a^{\prime},\omega\right)$ would mean that $\sigma$ cannot be S-BR. Given that any two messages sent in $\omega$ induce the same action, we can define $\rho\left(\sigma\left(\omega\right)\right)$ as the action that Receiver takes in state $\omega$ given $\left(\sigma,\rho\right)$.
[…]
It is perhaps worth noting that Theorems 1 and 2 jointly imply the following:
Corollary 1.
Generically, if cheap-talk Sender values randomization, then committed Sender values randomization.
5 Application to grading
For an application of our results, we consider their implications for grading policies. This application also clarifies a sense in which “randomization” in the statement of our results need not be interpreted literally.
Suppose Sender is a school that assigns grades to its students. We interpret $M$ as the set of potential grades. Each student is characterized by a vector of attributes. We say an attribute is relevant if an employer values it or the school’s value of placing the student with an employer depends on it. We interpret $\Omega$ as the set of all possible configurations of the relevant attributes. We maintain the assumption that $\Omega$ is finite.
Students also have irrelevant attributes. We denote by $X$ as the set of all possible configurations of the irrelevant attributes. We assume that the distribution over $X$ is atomless. The school utilizes a deterministic grading scheme $g:\Omega\times X\rightarrow M$. We say a grading scheme $g$ is fair if $g\left(\omega,x\right)=g\left(\omega,x^{\prime}\right)$ for every $\omega,x,x^{\prime}$. Otherwise, the scheme is unfair.
For this application, instead of envisioning a single Receiver, we assume that each student applies to a distinct employer. Each employer observes the grade $m\in M$ of its applicant and chooses one of finitely many actions $a\in A$ (e.g., whether to hire the student and if so for what position). All employers have the same utility function $u_{R}\left(a,\omega\right)$ that depends on the employer’s action and the relevant attributes of the applicant. (If there were a single employer who observed the grades of all of the applicants, this would effectively provide Sender with some commitment power because the distribution of messages would be directly observable to Receiver.) The school’s utility is additive across its students; for each student, the school’s payoff $u_{S}\left(a,\omega\right)$ depends on that student’s outcome and that student’s relevant attributes.
Under discretionary grading, the school freely chooses a grade to assign to each student, i.e., the school selects any grading scheme it wishes. The employer only observes its applicant’s grade but not the grading scheme that was used.
Alternatively, the school could implement a (publicly observable) grading policy that restricts the set of schemes that it can use. A grading policy could be a restriction to one specific grading scheme. This would make the situation equivalent to Bayesian persuasion. This is the case even though the grading scheme is deterministic because, by conditioning the grade on the irrelevant attributes, the school can implement any distribution of grades conditional on each $\omega$. [^18] The formulation of experiments as deterministic functions of an expanded state space was introduced by gentzkow2017bayesian and green2022two. It has been further studied in brooks2022information and brooks2024representing.
Another type of grading policy is one where the school commits to a given distribution of grades (lin2024credible). We refer to such a policy as a mandated curve. For example, the University of Chicago Law School mandates a pre-specified share of students that will receive a given narrow range of numerical grades.
More common is commitment to a GPA cap. For example, the University of Chicago Booth School of Business mandates that the average grade assigned in a given course must not exceed B+.
We say that the school values commitment if it strictly prefers to implement any grading policy (full commitment, mandated curve, GPA cap, etc.) over discretionary grading. We know that any policy must yield a payoff that is weakly lower than full commitment and weakly higher than discretionary grading. Consequently, if any grading policy yields a strictly higher payoff than discretionary grades, we know that the persuasion payoff (full commitment) exceeds the cheap talk payoff (discretionary grades).
We say that the school prefers to grade unfairly if its optimal grading scheme is unfair. In other words, if the school were able to commit to a particular grading scheme, it would select an unfair one.
Theorem 1 tells us that, generically, the school values commitment if and only if it prefers to grade unfairly. Thus, whenever we observe a school mandating a curve or a GPA cap, we know that the school’s ideal policy is unfair. [^19] Our analysis views the school (that cares about student placements) and the professor (who is assigning grades) as a single agent. A distinct motivation for a grading policy such as a GPA cap, outside of our Sender-Receiver framework, is an agency conflict between the school and the professor. For example, the professor may wish to give uniformly high grades in order to avoid student complaints so the school might impose a GPA cap to mitigate that temptation (frankel2014aligned). Moreover, a grading policy could have a distinct benefit of aiding equilibrium coordination about the meaning of grades; our focus on Sender-preferred equilibria assumes miscoordination away. Finally, our analysis takes the distribution of relevant attributes as exogenous. In practice, grading schemes not only provide information about the students but also incentivize the students to learn the material (boleslavsky2015grading).
Note, however, that even if we observe a school mandating a curve, Theorem 1 does not imply that the school will implement an unfair scheme if it can only commit to a mandated curve (i.e., is unable to fully commit to a particular scheme). In Online Appendix B.3, we analyze whether partial commitment being valuable (i.e., mandating a curve yields a strictly higher payoff than discretionary grades) implies that randomization under partial commitment is valuable (i.e., among the schemes that yield the mandated curve, every scheme that is optimal is unfair). Under the assumption that the school’s preferences are supermodular, we establish that this is indeed the case (Theorem 6). Whether the conclusion of this result holds when preferences are not supermodular remains an open question.
6 How often is commitment valuable?
Theorems 1 and 2 by themselves do not shed light on the types of environments where commitment (and thus randomization) are valuable nor about how common such environments are. This section aims to fill that gap.
Say that the environment is felicitous if Sender obtains his ideal payoff under cheap talk. [^20] We thank Roger Myerson for suggesting this term. It is immediate that commitment has no value if the environment is felicitous. The converse, of course, is not generally true (e.g, $u_{S}=-u_{R}$).
We first show that, for any $\Omega$, the share of environments that are felicitous is at least $\frac{1}{\left|A\right|^{\left|A\right|}}$. Thus, in general, the share of environments such that commitment has no value must be greater than $\frac{1}{\left|A\right|^{\left|A\right|}}$.
We then establish that, in the limit as $\Omega$ grows large: ($\ast$) the share of environments that are felicitous converges to exactly $\frac{1}{\left|A\right|^{\left|A\right|}}$, and ($\ast\ast$) generically, commitment has no value if and only if the environment is felicitous, i.e., the non-felicitous environments where commitment has no value (such as $u_{S}=-u_{R}$) become vanishingly rare. Thus, as $\left|\Omega\right|\rightarrow\infty$, the share of environments such that commitment has no value converges exactly to $\frac{1}{\left|A\right|^{\left|A\right|}}$.
Theorem 3.
Fix $A$.
  1. (i)
    For any $\Omega$, the share of environments such that commitment has no value is at least $\frac{1}{\left|A\right|^{\left|A\right|}}$ .
  2. (ii)
    As $\left|\Omega\right|\rightarrow\infty$,
    1. (a)
      generically, commitment has no value if and only if Sender obtains his ideal payoff under cheap talk.
    2. (b)
      the share of environments such that commitment has no value converges to $\frac{1}{\left|A\right|^{\left|A\right|}}$.
Here we provide a sketch of the proof. For each $a\in A$, let $\Omega_{a}$ be the set of states where $a$ is Sender’s ideal action. Generically, if $a\neq a^{\prime}$, $\Omega_{a}$ and $\Omega_{a^{\prime}}$ do not intersect; consider this case. Then, it is easy to see that the environment is felicitous if and only if for each $a,a^{\prime}\in A$, we have [^21] This condition also appears in antic2022subversive and aybas2024cheap. In antic2022subversive, it is a necessary condition for the possibility of subversive conversations: without it, a third-party with veto power would prevent a committee from implementing a project solely based on the information that the committee wants to do so. aybas2024cheap consider preferences of the form $u_{R}\left(a,\omega\left(\cdot\right)\right)=\omega\left(a\right)^{2}$ and $u_{S}\left(a,\omega\left(\cdot\right)\right)=\left(\omega\left(a\right)-b\right)^{2}$ for some $b>0$ where $\omega:A\rightarrow\mathbb{R}$ is the realized path of a Brownian motion. They identify features of $b$ and $A$ such that the condition holds.
$\sum_{\omega\in\Omega_{a}}\mu_{0}\left(\omega\right)\left(u_{R}\left(a,\omega\right)-u_{R}\left(a^{\prime},\omega\right)\right)\geq 0.$ (1)
[…]
We establish part (ii) of the theorem by showing that as $\left|\Omega\right|$ grows large: ($\ast$) the share of preference such that an $\Omega_{a}$ is empty converges to zero so the share of environments that are felicitous converges to $\frac{1}{\left|A\right|^{\left|A\right|}}$, and ($\ast\ast$) the share of environments such that commitment has no value converges to the share of environments that are felicitous.
Part ($\ast$) is easy to see. For any $a\in A$, as $\Omega$ grows large, the share of preferences such that there is no state where $a$ is Sender’s ideal action converges to zero.
To establish part ($\ast\ast$), say that an environment is jointly-inclusive if for every action $a$, there is some state $\omega$ such that $a$ is the ideal action for both Sender and Receiver in $\omega$. Analogously to part ($\ast$), it is easy to see that as $\Omega$ grows large, the share of environments that are jointly-inclusive converges to $1$. To complete the proof of part ($\ast\ast$), we argue that, generically, if the environment is jointly-inclusive and commitment has no value, then the environment must be felicitous. First, because commitment has no value, Theorem 2 implies that there is a partitional profile $\left(\sigma,\rho\right)$ that is a cheap-talk equilibrium and yields the persuasion payoff. Next, we note that every action $a\in A$ must be induced by $\left(\sigma,\rho\right)$: there is a state $\omega$ where $a$ is both Sender’s and Receiver’s ideal action, so if $a$ were never taken, the committed Sender could profitably deviate by sometimes [^22] Sender could reveal $\omega$ with some probability $\varepsilon$; Receiver’s response to all other messages would remain unchanged if $\varepsilon$ is sufficiently small. revealing $\omega$ and inducing $a$, thus contradicting the fact that $\left(\sigma,\rho\right)$ yields the persuasion payoff. This in turn implies that, for every $\omega$, $\rho\left(\sigma\left(\omega\right)\right)$ must be Sender’s ideal action in $\omega$. (If Sender strictly preferred some other $a^{\prime}$ in $\omega$, $\left(\sigma,\rho\right)$ could not be S-BR as the cheap-talk Sender would profitably deviate and set $\sigma$$\left(\omega\right)$ to be whatever message induces $a^{\prime}$; since all actions are induced by $\left(\sigma,\rho\right)$, there must be such a message.) Thus, $\left(\sigma,\rho\right)$ is a partitional profile that is R-BR and induces Receiver to take Sender’s ideal action in every state. But this means that every message sent under $\sigma$ fully reveals what action is ideal for Sender, and Receiver complies and takes that action. Hence, the environment is felicitous.
We conclude this section with a few comments.
First, whether commitment has value in a given environment $\left(u_{S},u_{R}\right)$ depends on the prior $\mu_{0}$. Yet, Theorem 3 remarkably holds for any (interior) prior.
Second, a corollary of Theorem 3, part 2b, is that in the limit $\lim_{\left|A\right|\rightarrow\infty}\lim_{\left|\Omega\right|\rightarrow\infty}$, commitment generically has value. One might be tempted to summarize this corollary as “if both $A$ and $\Omega$ are large, then commitment is almost certainly valuable.” An issue with this summary, however, is that, at least given our current proof method, the conclusion relies on the order of limits. Whether commitment is generically valuable when the order of the limits is reversed (i.e., $\lim_{\left|\Omega\right|\rightarrow\infty}\lim_{\left|A\right|\rightarrow\infty})$ or we consider the joint limit (i.e., $\lim_{\begin{subarray}{c}|\Omega|\to\infty\\ |A|\to\infty\end{subarray}}$) remain open questions.
Third, the felicity condition has a flavor of alignment of Sender and Receiver’s preferences. However, it does not preclude the possibility that Receiver is much worse off than she would be if Sender and Receiver’s preferences were fully aligned. For instance, consider the prosecutor-judge example and suppose that the prior is $0.7$ rather than $0.3$; then, the environment is felicitous but Receiver obtains no information.
Appendix A Appendix
A.1 Notation and terminology
Let $A=\left\{a_{1},...,a_{\left|A\right|}\right\}$. Let $\Omega=\left\{\omega_{1},...,\omega_{\left|\Omega\right|}\right\}$.
Given a messaging strategy $\sigma$, let $M_{\sigma}=\{m\in M|\sigma(m|\omega)>0\text{ for some }\omega\}$ be the set of messages that are sent with positive probability under $\sigma$. For any $\omega$, if $\sigma(\cdot|\omega)$ is degenerate (i.e., there exists a message $m$ such that $\sigma(m|\omega)=1$), let $\sigma(\omega)$ denote the message that is sent in state $\omega$. Similarly, if $\rho(\cdot|m)$ is degenerate, let $\rho\left(m\right)$ denote the action taken following message $m$.
Say that $\rho$ is pure if $\rho(\cdot|m)$ is degenerate for all $m\in M$. Given a profile $\left(\sigma,\rho\right)$, say $\rho$ is pure-on-path if $\rho\left(\cdot|m\right)$ is degenerate for all $m\in M_{\sigma}$.
Let $I=\left[\begin{array}[]{c}e_{1}\\ ...\\ e_{\left|\Omega\right|}\end{array}\right]$ be the identity matrix of size $\left|\Omega\right|$, with $e_{i}$ being the row vector with all entries equal to 0 except for the $i^{th}$ entry equal to $1$. Denote a vector all of whose elements are equal to $r$ by $\boldsymbol{r}$. Denote the $j^{th}$ element of $\mu$ by $\left[\mu\right]_{j}$.
A.2 Generic environments for the proofs
We now introduce two generic sets of environments that play important roles in the proofs.
A.2.1 Partitional-unique-response environments
An environment $\left(u_{S},u_{R}\right)$ satisfies partitional-unique-response if for every non-empty $\hat{\Omega}\subseteq\Omega$,
$\arg\max_{a\in A}\sum_{\omega\in\hat{\Omega}}\mu_{0}(\omega)u_{R}(a,\omega)$
is a singleton.
Note that whether an environment satisfies partitional-unique-response does not depend on Sender’s preferences. The partitional-unique-response property requires that, at the finitely many beliefs that can be induced by a partitional experiment, Receiver has a unique best response.
Lemma 1.
The set of partitional-unique-response environments is generic.
Proof.
Let $\mathcal{T}$ denote the set of triplets $\left(\hat{\Omega},a_{i},a_{j}\right)$ such that $\emptyset\neq\hat{\Omega}\subseteq\Omega$, $a_{i},a_{j}\in A$, and $a_{i}\neq a_{j}$. For any $\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}$, let $Q(\hat{\Omega},a_{i},a_{j})$ denote the set of $u_{R}$ such that
$\sum_{\omega\in\hat{\Omega}}\,\mu_{0}(\omega)u_{R}(a_{i},\omega)=\sum_{\omega\in\hat{\Omega}}\,\mu_{0}(\omega)u_{R}(a_{j},\omega).$ (2)
We will show that $\cup_{\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}}Q(\hat{\Omega},a_{i},a_{j})$ has measure zero in $0,1]^{|A||\Omega|}$. Since $A$ and $\Omega$ are finite, it suffices to show that for any $\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}$, $Q(\hat{\Omega},a_{i},a_{j})$ has measure zero.
Fix any $\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}$. Note that $Q(\hat{\Omega},a_{i},a_{j})$ can be written as
$\left\{u_{R}\in\left[0,1\right]^{|A||\Omega|}|\sum_{\omega,a}u_{R}(a,\omega)\eta(a,\omega)=0\right\}$ (3)
where
$\eta(a,\omega)=\begin{cases}\mu_{0}(\omega)&\text{ if }a=a_{i},\omega\in\hat{\Omega}\\ -\mu_{0}(\omega)&\text{ if }a=a_{j},\omega\in\hat{\Omega}\\ 0&\text{ otherwise.}\end{cases}$
Hence, $Q(\hat{\Omega},a_{i},a_{j})$ is a subset of a hyperplane in $\mathbb{R}^{|A||\Omega|}$, and thus has measure zero. [^23] The set of partitional-unique-response environment is also generic in the topological sense. It is easy to see that $Q(\hat{\Omega},a_{i},a_{j})$ is closed. Since $\cup_{\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}}Q(\hat{\Omega},a_{i},a_{j})$ is therefore closed, its complement is open. Since $\cup_{\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}}Q(\hat{\Omega},a_{i},a_{j})$ has measure zero, its complement is dense. Thus, the set of partitional-unique-response environments, which is a superset of the complement of $\cup_{\left(\hat{\Omega},a_{i},a_{j}\right)\in\mathcal{T}}Q(\hat{\Omega},a_{i},a_{j})$, contains an open, dense set.
A.2.2 Scant-indifferences environments
For each $a_{i}\in A$, let $\mathbf{u}_{S}(a_{i})=u_{S}(a_{i},\cdot)\in\mathbb{R}^{\left|\Omega\right|}$ and $\mathbf{u}_{R}(a_{i})=u_{R}(a_{i},\cdot)\in\mathbb{R}^{|\Omega|}$ denote the payoff vectors across states.
For each $a_{i},$ define the expanded-indifference matrix $T^{i}$ as follows. Let $T_{S}^{i}$ be the matrix with $\left|A\right|-1$ rows and $\left|\Omega\right|$ columns, with each row associated with $j\neq i$ and equal to $\mathbf{u}_{S}(a_{j})-\mathbf{u}_{S}(a_{i})$. Let $T_{R}^{i}$ be the matrix with $\left|A\right|-1$ rows and $\left|\Omega\right|$ columns, with each row associated with $j\neq i$ and equal to $\mathbf{u}_{R}(a_{j})-\mathbf{u}_{R}(a_{i})$. Then, let
$T^{i}=\begin{bmatrix}T_{S}^{i}\\ T_{R}^{i}\\ I\end{bmatrix}$.
Given any matrix $T$, a row-submatrix of $T$ is a matrix formed by removing some of the rows of $T$.
We say that an environment satisfies scant-indifferences if for each $a_{i}\in A$, every row-submatrix of the expanded-indifference matrix $T^{i}$ is full rank.
We anticipate that the reader might find this definition mysterious, so we now try to provide some intuition by connecting this definition to the proof sketch we gave in the body of the paper for Theorem 2 in the case with two actions and three states.
Recall, that in Figure 1, the argument behind Lemma 4 relied on two facts that must hold generically. First, the border between $R_{1}$ and $R_{2}$ is distinct from the border between $S_{1}$ and $S_{2}$ and thus the two borders have at most one intersection, $\mu_{m}$. Second, generically $\mu_{m}$ (if it exists) is an interior belief. Moreover, the argument behind Lemma 5 relied on the fact that, generically, for any $\omega$ and $a_{i}\neq a_{j}$, $u_{S}(a_{i},\omega)\neq u_{S}(a_{j},\omega)$.
We now illustrate why these three facts hold in any scant-indifferences environment. With only two actions, we can look at $T^{1}$ only, since the argument for $T^{2}$ is identical. We have
$T^{1}=\begin{bmatrix}\begin{array}[]{ccc}\overset{\Delta}{u}_{S}\left(\omega_{1}\right)&\overset{\Delta}{u}_{S}\left(\omega_{2}\right)&\overset{\Delta}{u}_{S}\left(\omega_{3}\right)\\ \overset{\Delta}{u}_{R}\left(\omega_{1}\right)&\overset{\Delta}{u}_{R}\left(\omega_{2}\right)&\overset{\Delta}{u}_{R}\left(\omega_{3}\right)\\ 1&0&0\\ 0&1&0\\ 0&0&1\end{array}\end{bmatrix}$
where $\overset{\Delta}{u}_{S}\left(\omega_{i}\right)=u_{S}\left(a_{2},\omega_{i}\right)-u_{S}\left(a_{1},\omega_{i}\right)$ and analogously for $\overset{\Delta}{u}_{R}$.
First, consider the row-submatrix
$\overset{\Delta}{T}=\begin{bmatrix}\overset{\Delta}{u}_{S}\left(\omega_{1}\right)&\overset{\Delta}{u}_{S}\left(\omega_{2}\right)&\overset{\Delta}{u}_{S}\left(\omega_{3}\right)\\ \overset{\Delta}{u}_{R}\left(\omega_{1}\right)&\overset{\Delta}{u}_{R}\left(\omega_{2}\right)&\overset{\Delta}{u}_{R}\left(\omega_{3}\right)\end{bmatrix}$.
Note that both Sender and Receiver are indifferent between the two actions at a belief $\mu$ if and only if $\overset{\Delta}{T}\mu=0$. Thus, requiring that $\overset{\Delta}{T}$ be full-rank is equivalent to requiring that the border between $R_{1}$ and $R_{2}$ not be parallel to the border between $S_{1}$ and $S_{2}$. A fortiori, the environment satisfying scant-indifferences implies that the two borders do not coincide.
Second, consider the row-submatrix
$\begin{bmatrix}\overset{\Delta}{u}_{S}\left(\omega_{1}\right)&\overset{\Delta}{u}_{S}\left(\omega_{2}\right)&\overset{\Delta}{u}_{S}\left(\omega_{3}\right)\\ \overset{\Delta}{u}_{R}\left(\omega_{1}\right)&\overset{\Delta}{u}_{R}\left(\omega_{2}\right)&\overset{\Delta}{u}_{R}\left(\omega_{3}\right)\\ 1&0&0\end{bmatrix}$.
Requiring that this matrix be full-rank yields that $\mu_{m}$ puts strictly positive probability on $\omega_{1}$. Considering the row-submatrices that alternatively include the other two rows of the identity matrix yields that $\mu_{m}$ puts strictly positive probability on $\omega_{2}$ and $\omega_{3}$.
Finally, suppose that in, say state $\omega_{1}$, $\overset{\Delta}{u}_{S}\left(\omega_{1}\right)=0$. Consider the row-submatrix
$\begin{bmatrix}0&\overset{\Delta}{u}_{S}\left(\omega_{2}\right)&\overset{\Delta}{u}_{S}\left(\omega_{3}\right)\\ 0&1&0\\ 0&0&1\end{bmatrix}$.
Clearly, this matrix is not full-rank, so scant-indifferences rules out the possibility that $u_{S}(a_{1},\omega_{1})=u_{S}(a_{2},\omega_{1})$.
Having motivated the definition of scant-indifferences environments, we now establish that the set of such environments is generic.
Lemma 2.
The set of scant-indifferences environments is generic.
Proof.
First, observe that given any expanded-indifference matrix $T^{i}$, if every square row-submatrix of $T^{i}$ is full-rank, then every row-submatrix of $T^{i}$ is full-rank. To see why, suppose every square row-submatrix of $T^{i}$ is full-rank. Now, consider an arbitrary row-submatrix $\hat{T}$ of $T^{i}$. If $\hat{T}$ is square, it obviously has full-rank. Suppose that $\hat{T}$ has more than $\left|\Omega\right|$ rows. In that case, every square row-submatrix of $\hat{T}$ is also a square row-submatrix of $T^{i}$. This row-submatrix has rank $\left|\Omega\right|$. Therefore, $\hat{T}$ has rank $\left|\Omega\right|$ and is thus full-rank. Finally, suppose that $\hat{T}$ has fewer than $\left|\Omega\right|$ rows. We know that $\hat{T}$ is a row-submatrix of some square row-submatrix $\tilde{T}$ of $T^{i}$. We know $\tilde{T}$ has full-rank so all of its rows are linearly independent. Consequently, the subset of its rows that constitute $\hat{T}$ is also linearly independent.
Hence, we can consider only square row-submatrices of $T^{i}$. Recall that a square matrix is full-rank if and only if its determinant is non-zero. Thus, it will suffice to show that for a full Lebesgue measure set of $\left(u_{S},u_{R}\right)$, the determinant of every square row-submatrix of each expanded-indifference matrix is non-zero. Consider any square row-submatrix $\hat{T}$ of any expanded-indifference matrix. The determinant of $\hat{T}$ is a non-zero polynomial function of $(u_{S},u_{R})$. [^24] To establish that this polynomial function is not identically zero, we need to show that there are values of $(u_{S},u_{R})$ such that the determinant of $\hat{T}$ is not zero. Recall that rearranging the rows of a matrix does not change whether the determinant is zero. Let $\tilde{T}$ be some matrix constructed by rearranging the rows of $\hat{T}$ so that if $\tilde{T}$ contains any row equal to $e_{k}$, that row is the $k^{th}$ row of $\tilde{T}$. Any other row $\ell$ of $\tilde{T}$ must belong either to $T_{S}^{i}$ or $T_{R}^{i}$. Construct $\left(u_{S},u_{R}\right)$ as follows. If $\mathbf{u}_{S}(a_{j})-\mathbf{u}_{S}(a_{i})$ is not in $\hat{T}$ (and thus not in $\tilde{T}$), set $\mathbf{u}_{S}(a_{j})$ in an arbitrary way. If $\mathbf{u}_{S}(a_{j})-\mathbf{u}_{S}(a_{i})$ is the $\ell^{th}$ row of $\tilde{T}$, set $\mathbf{u}_{S}(a_{j})$ as the row vector with all entries equal to 0 except for the $\ell^{th}$ entry equal to $1$. Similarly, if $\mathbf{u}_{R}(a_{j})-\mathbf{u}_{R}(a_{i})$ is not in $\hat{T}$ (and thus not in $\tilde{T}$), set $\mathbf{u}_{R}(a_{j})$ in an arbitrary way, and if $\mathbf{u}_{R}(a_{j})-\mathbf{u}_{R}(a_{i})$ is the $\ell^{th}$ row of $\tilde{T}$, set $\mathbf{u}_{R}(a_{j})$ as the row vector with all entries equal to 0 except for the $\ell^{th}$ entry equal to $1$. Finally, set $\mathbf{u}_{S}(a_{i})=\mathbf{u}_{R}(a_{i})=\boldsymbol{0}$. With $(u_{S},u_{R})$ defined this way, $\tilde{T}$ is an identity matrix with determinant 1. Therefore, the determinant of $\hat{T}$ is a non-zero function of $(u_{S},u_{R})$. The zero set of any non-zero polynomial function has Lebesgue measure zero, so the set of $\left(u_{S},u_{R}\right)$ for which $\hat{T}$ does not have full rank is a measure-zero set. Since there are only finitely many square row-submatrices of expanded-indifference matrices, the set of scant-indifferences environments is generic. [^25] The zero set of any non-zero polynomial function is closed, so the set of scant indifferences environments is generic in the topological sense as well.
As we noted above (for the three state, two action case), in scant-indifferences environments, there is no state in which Sender is indifferent between two distinct actions.
Lemma 3.
In any scant-indifferences environment, for any $\omega$ and $a_{i}\neq a_{j}$, $u_{S}(a_{i},\omega)\neq u_{S}(a_{j},\omega)$.
Proof.
Suppose, toward a contradiction, that there exist some $\omega$, $a_{i}$, and $a_{j}$ such that $u_{S}(a_{i},\omega)=u_{S}(a_{j},\omega)$. Without loss, suppose this holds for $\omega_{1}$. Then, the vector $\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{j})$ has zero as its first element. Now consider the $|\Omega|\times|\Omega|$ row sub-matrix of $T^{j}$
$\begin{bmatrix}\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{j})\\ e_{2}\\ ...\\ e_{|\Omega|}\end{bmatrix}$.
This matrix is not full-rank because the first row can be expressed as a linear combination of the other rows. ∎
A.3 Key Proposition
In this section we establish a key proposition that will be useful for proofs of Theorems 1, 2, and 3.
Proposition 1.
[…]
To establish the Proposition, we first show that if a cheap-talk equilibrium (in fact any R-BR profile) yields the persuasion payoff, then Receiver must not randomize on path in that equilibrium (Lemma 4). Second, we show that if Receiver does not randomize on path, Sender also need not randomize (Lemma 5).
Lemma 4.
In a scant-indifferences environment, if $\left(\sigma,\rho\right)$ is R-BR and yields the persuasion payoff, then $\rho$ must be pure-on-path.
Proof.
Suppose by contradiction that the environment satisfies scant-indifferences, profile $\left(\sigma,\rho\right)$ is R-BR and yields the persuasion payoff, yet there exists a message $m\in M_{\sigma}$ such that $|\operatorname{Supp}(\rho(\cdot|m))|=k>1$.
We first note that both Sender and Receiver must be indifferent among all the actions in $\operatorname{Supp}(\rho(\cdot|m))$ given $\mu_{m}$, the belief induced by message $m$. In other words, for all $a_{i},a_{j}\in\operatorname{Supp}(\rho(\cdot|m))$,
$\sum_{\omega}\mu_{m}(\omega)u_{R}(a_{i},\omega)=\sum_{\omega}\mu_{m}(\omega)u_{R}(a_{j},\omega),$ (4)
$\sum_{\omega}\mu_{m}(\omega)u_{S}(a_{i},\omega)=\sum_{\omega}\mu_{m}(\omega)u_{S}(a_{j},\omega).$ (5)
Equation (4) follows immediately from R-BR. Equation (5) follows from the fact that $(\sigma,\rho)$ yields the persuasion payoff: if say $\sum_{\omega}\mu_{m}(\omega)u_{S}(a_{i},\omega)>\sum_{\omega}\mu_{m}(\omega)u_{S}(a_{j},\omega)$, an alternative strategy profile where Receiver breaks ties in favor of Sender would still satisfy R-BR while strictly improving Sender’s payoff.
For each belief $\mu\in\Delta\Omega$, let $A_{R}^{*}(\mu)$ denote the set of Receiver-optimal actions under belief $\mu$; that is, $A_{R}^{*}(\mu)=\arg\max_{a\in A}\mathbf{u}_{R}(a)\cdot\mu$. Clearly, $\operatorname{Supp}(\rho(\cdot|m))\subseteq A_{R}^{*}(\mu_{m})$, meaning that $A_{R}^{*}(\mu_{m})$ contains the $k$ actions in the support of $\rho(\cdot|m)$, but may also contain additional actions that are not played following $m$. Without loss of generality, let $\operatorname{Supp}(\rho(\cdot|m))=\{a_{1},...,a_{k}\}$ and $A_{R}^{*}(\mu_{m})=\{a_{1},...,a_{k},a_{k+1},...,a_{k+r}\}$ for some $r\geq 0$. Note that for any i=2,,k+r,i=2,...,k+r, $\mathbf{u}_{R}(a_{1})\cdot\mu_{m}=\mathbf{u}_{R}(a_{i})\cdot\mu_{m}$.
Equation (5) implies that for any i=2,,k,i=2,...,k, $\mathbf{u}_{S}(a_{1})\cdot\mu_{m}=\mathbf{u}_{S}(a_{i})\cdot\mu_{m}$. Combining both Sender’s and Receiver’s indifference conditions, we have
$\begin{bmatrix}\mathbf{u}_{S}(a_{2})-\mathbf{u}_{S}(a_{1})\\ ...\\ \mathbf{u}_{S}(a_{k})-\mathbf{u}_{S}(a_{1})\\ \\ \mathbf{u}_{R}(a_{2})-\mathbf{u}_{R}(a_{1})\\ ...\\ \mathbf{u}_{R}(a_{k+r})-\mathbf{u}_{R}(a_{1})\end{bmatrix}\mu_{m}=\boldsymbol{0}.$ (6)
Let $\hat{\Omega}=\{\omega|\mu_{m}(\omega)=0\}$, the (potentially empty) set of states that are not in the support of $\mu_{m}$. Without loss, suppose that $\hat{\Omega}=\left\{\omega_{1},...\omega_{\ell}\right\}$ where $\ell\geq 0$. If $\ell>0$ (i.e., $\hat{\Omega}\neq\emptyset$), then we have
$\begin{bmatrix}e_{1}\\ ...\\ e_{\ell}\end{bmatrix}\mu_{m}=\boldsymbol{0}.$ (7)
Let $\hat{T}_{S}=\begin{bmatrix}\mathbf{u}_{S}(a_{2})-\mathbf{u}_{S}(a_{1})\\ ...\\ \mathbf{u}_{S}(a_{k})-\mathbf{u}_{S}(a_{1})\end{bmatrix}$, $\hat{T}_{R}=\begin{bmatrix}\mathbf{u}_{R}(a_{2})-\mathbf{u}_{R}(a_{1})\\ ...\\ \mathbf{u}_{R}(a_{k+r})-\mathbf{u}_{R}(a_{1})\end{bmatrix}$, $\hat{E}=\begin{bmatrix}e_{1}\\ ...\\ e_{\ell}\end{bmatrix}$, and $\hat{T}=\begin{bmatrix}\hat{T}_{S}\\ \hat{T}_{R}\\ \hat{E}\end{bmatrix}$. Note that $\hat{T}$ is a row-submatrix of the expanded-indifference matrix $T^{1}$.
Combining (6) and (7), we know $\hat{T}\mu_{m}=\boldsymbol{0}$. Moreover, since $\mu_{m}\in\Delta\Omega$, we know $\boldsymbol{1}\mu_{m}=1$.
Next we make two observations: (i) $rank(\hat{T})<|\Omega|$, otherwise the unique solution to $\text{$\hat{T}$}\mu=\boldsymbol{0}$ is $\mu=\boldsymbol{0}$. Since we are in a scant-indifferences environment, this means that $\hat{T}$ has full row rank; (ii) vector $\boldsymbol{1}$ can not be represented as a linear combination of rows of $\hat{T}$. To see why, assume toward contradiction that there exists a row vector $\lambda\in\mathbb{R}^{2k+r+\ell-2}$ such that $\lambda\hat{T}=\boldsymbol{1}$. This would lead to a contradiction that $1=\boldsymbol{1}\mu_{m}=\lambda\hat{T}\mu_{m}=\lambda\boldsymbol{0}=0$.
Observations (i) and (ii) together imply that the matrix $\begin{bmatrix}\hat{T}\\ \boldsymbol{1}\end{bmatrix}$ has full row rank. Consequently, we know $rank\left(\left[\begin{array}[]{c}\text{$\hat{T}$}\\ \boldsymbol{1}\end{array}\right]\right)>rank\begin{pmatrix}\begin{bmatrix}\hat{T}_{R}\\ \hat{E}\\ \boldsymbol{1}\end{bmatrix}\end{pmatrix}$.
Now, we claim that there exists $x\in\mathbb{R}^{n}$ such that
$\begin{bmatrix}\hat{T}_{R}\\ \hat{E}\\ \boldsymbol{1}\end{bmatrix}x=0$ (8)
and
$\hat{T}_{S}\,x\neq 0.$ (9)
To see this, suppose by contradiction that for any $x$ that solves (8), we have $\hat{T}_{S}\,x=0$. This would imply that the set of solutions to (8) and the set of solutions to
$\begin{bmatrix}\hat{T}\\ \boldsymbol{1}\end{bmatrix}x=0$ (10)
[…]
Consider an $x$ that satisfies (8) and (9). Consider two vectors, $\mu_{m}+\varepsilon x$ and $\mu_{m}-\varepsilon x$, for some $\varepsilon\in\mathbb{R}_{>0}$. First we verify that for sufficiently small $\varepsilon$, $\mu_{m}\pm\varepsilon x\in\Delta\text{$\Omega$}$. Since $\boldsymbol{1}x=0$, it follows that $\boldsymbol{1}\left(\mu_{m}\pm\varepsilon x\right)=\boldsymbol{1}\mu_{m}=1$. For $\omega_{j}\notin\hat{\Omega}$, we have $\left[\mu_{m}\right]_{j}>0$, so for small enough $\varepsilon$, $\left[\mu_{m}\pm\varepsilon x\right]_{j}\geq 0$. For $\omega_{j}\in\hat{\Omega}$, we know $e_{j}$ is a row of $\hat{E}$, so $e_{j}x=0$. Consequently, $\left[\mu_{m}\pm\varepsilon x\right]_{j}=e_{j}\left(\mu_{m}\pm\varepsilon x\right)=\left[\mu_{m}\right]_{j}=0.$ Thus, $\mu_{m}\pm\varepsilon x\in\Delta\text{$\Omega$}$.
Observe that $A_{R}^{*}(\mu_{m})=A_{R}^{*}(\mu_{m}\pm\varepsilon x)$. First, for any $a\notin A_{R}^{*}(\mu_{m})$, if $\varepsilon$ is sufficiently small, $a\notin A_{R}^{*}(\mu_{m}\pm\varepsilon x)$. Therefore, $A_{R}^{*}(\mu_{m}\pm\varepsilon x)\subseteq A_{R}^{*}(\mu_{m})$. But, $\hat{T}_{R}\,x=0$ implies that $\left(\mu_{m}\pm\varepsilon x\right)\cdot\mathbf{u}_{R}(a)$ is constant across $a\in A_{R}^{*}(\mu_{m})$, so $A_{R}^{*}(\mu_{m}\pm\varepsilon x)=A_{R}^{*}(\mu_{m})$.
[…]
Since $\hat{T}_{S}\,x\neq 0$, we know there is an $a_{i}\in\{a_{2},...,a_{k}\}$ such that $x\cdot\left(\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{1})\right)\neq 0$.
Because $a_{1}\in A_{R}^{*}(\mu_{m}\pm\varepsilon x)=A_{R}^{*}(\mu_{m})$, we have
$\max_{a\in A^{*}(\mu_{m})}\left(\mu_{m}+\varepsilon x\right)\cdot\left(\mathbf{u}_{S}(a)-\mathbf{u}_{S}(a_{1})\right)\geq 0$
and
$\max_{a\in A^{*}(\mu_{m})}\left(\mu_{m}-\varepsilon x\right)\cdot\left(\mathbf{u}_{S}(a)-\mathbf{u}_{S}(a_{1})\right)\geq 0.$
We now establish that at least one of these inequalities has to be strict. Suppose toward contradiction that both hold with equality. The first equality implies that for all $a_{i}\in\{a_{2},...,a_{k}\}$, $\left(\mu_{m}+\varepsilon x\right)\cdot\left(\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{1})\right)\leq 0$, which combined with the fact that $\mu_{m}\cdot\mathbf{u}_{S}(a_{i})=\mu_{m}\cdot\mathbf{u}_{S}(a_{1})$ implies that $x\cdot\left(\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{1})\right)\leq 0$ for all $a_{i}\in\{a_{2},...,a_{k}\}$. Similarly, the second equality implies that $x\cdot\left(\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{1})\right)\leq 0$ for all $a_{i}\in\{a_{2},...,a_{k}\}$. Together, this yields that $x\cdot\left(\mathbf{u}_{S}(a_{i})-\mathbf{u}_{S}(a_{1})\right)=0$ for all $a_{i}\in\{a_{2},...,a_{k}\}$, a contradiction. Hence, one of the inequalities has to be strict.
Consequently, Sender’s interim payoff under $\hat{\sigma}$ (in the event that $m$ is sent under $\sigma$) is
$\frac{1}{2}\max_{a\in A^{*}(\mu_{m})}\left(\mu_{m}+\varepsilon x\right)\cdot\mathbf{u}_{S}(a)+\frac{1}{2}\max_{a\in A^{*}(\mu_{m})}\left(\mu_{m}-\varepsilon x\right)\cdot\mathbf{u}_{S}(a)$
$>$ $\frac{1}{2}\left(\mu_{m}+\varepsilon x\right)\cdot\mathbf{u}_{S}(a_{1})+\frac{1}{2}\left(\mu_{m}-\varepsilon x\right)\cdot\mathbf{u}_{S}(a_{1})$
$=$ $\mu_{m}\cdot\mathbf{u}_{S}(a_{1})$
[…]
Lemma 5.
[…]
Proof.
Suppose a cheap-talk equilibrium $(\sigma,\rho)$ yields the persuasion payoff and $\rho$ is pure-on-path.
First, we show that for any $\omega$ and any $m,m^{\prime}$ such that $\sigma(m|\omega),\sigma(m^{\prime}|\omega)>0$, $\rho(m)=\rho(m^{\prime})$. The fact that both $m$ and $m^{\prime}$ are sent in $\omega$ implies, by S-BR, that $u_{S}(\rho(m),\omega)=u_{S}(\rho(m^{\prime}),\omega)$. Moreover, by Lemma 3, there exist no distinct $a$ and $a^{\prime}$ such that $u_{S}(a,\omega)=u_{S}(a^{\prime},\omega)$, so it must be that $\rho(m)=\rho(m^{\prime})$.
Let $A^{*}=\{a\in A|a=\rho(m)\text{ for some }m\in M_{\sigma}\}$ be the set of actions that are taken on-path. Without loss, let $A^{*}=\{a_{1},...,a_{k}\}$. For each $a_{i}$, let $M_{i}=\{m\in M_{\sigma}|\rho(m)=a_{i}\}$ be the set of on-path messages that induce action $a_{i}$, and $\Omega_{i}=\{\omega\in\Omega|\operatorname{Supp}(\sigma(\cdot|\omega))\subseteq M_{i}\}$ be the set of states that induce action $a_{i}$. Note that $\left\{M_{i}\right\}_{i=1}^{k}$ is a partition of $M_{\sigma}.$ Moreover, it is easy to see that $\{\Omega_{i}\}_{i=1}^{k}$ is a partition of $\Omega$. First, $\Omega_{i}$ cannot be empty because every $a_{i}\in A^{*}$ is taken on-path. Second, every $\omega\in\Omega$ belongs to some $\Omega_{i}$ as only actions in $A^{*}$ are taken on-path; hence, $\cup_{i}\Omega_{i}=\Omega$. Finally, the fact that for any $\omega$ and any $m,m^{\prime}$ such that $\sigma(m|\omega),\sigma(m^{\prime}|\omega)>0$ we have $\rho(m)=\rho(m^{\prime})$ implies that if $i\neq j$, $\Omega_{i}$ and $\Omega_{j}$ are disjoint.
[…]
  • $\hat{\sigma}(m_{i}|\omega)=1$ if $\omega\in\Omega_{i}$.
  • $\hat{\rho}(m_{i})=a_{i}.$
  • $\hat{\rho}(m)=a_{1}$ if $m\in M\backslash\{m_{1},....m_{k}\}$.
[…]
Note that S-BR of $(\sigma,\rho)$ implies that for any $\omega$ and $m\in\operatorname{Supp}(\sigma(\cdot|\omega))$, we have
$u_{S}(\rho(m),\omega)\geq u_{S}(\rho(m^{\prime}),\omega)\text{ for all }m^{\prime}\in M_{\sigma}.$
[…]
The fact that $(\sigma,\rho)$ is R-BR implies that for all $m\in M_{\sigma}$,
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\sigma(m|\omega)u_{R}(\rho(m),\omega)\geq\sum_{\omega\in\Omega}\mu_{0}(\omega)\sigma(m|\omega)u_{R}(a^{\prime},\omega)\quad\text{for all }a^{\prime}\in A.$
For any $i\in\left\{1,...,k\right\}$, we can sum the inequality above over $m\in M_{i}$. Since for $m\in M_{i}$ we have $\rho\left(m\right)=a_{i}$, this yields
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\sum_{m\in M_{i}}\sigma(m|\omega)u_{R}(a_{i},\omega)\geq\sum_{\omega\in\Omega}\mu_{0}(\omega)\sum_{m\in M_{i}}\sigma(m|\omega)u_{R}(a^{\prime},\omega)\quad\text{for all }a^{\prime}\in A.$
Since for any $m\in M_{i}$ and $\omega\notin\Omega_{i}$, we have $\sigma(m|\omega)=0$, the inequality above implies
$\sum_{\omega\in\Omega_{i}}\mu_{0}(\omega)\sum_{m\in M_{i}}\sigma(m|\omega)u_{R}(a_{i},\omega)\geq\sum_{\omega\in\Omega_{i}}\mu_{0}(\omega)\sum_{m\in M_{i}}\sigma(m|\omega)u_{R}(a^{\prime},\omega)\quad\text{for all }a^{\prime}\in A.$
Since $\sum_{m\in M_{i}}\sigma(m|\omega)=1$ if $\omega\in\Omega_{i}$, we have
$\sum_{\omega\in\Omega_{i}}\mu_{0}(\omega)u_{R}(a_{i},\omega)\geq\sum_{\omega\in\Omega_{i}}\mu_{0}(\omega)u_{R}(a^{\prime},\omega)\quad\text{for all }a^{\prime}\in A.$ (11)
[…]
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\hat{\sigma}\left(m_{i}|\omega\right)\sum_{a\in A}\hat{\rho}\left(a|m_{i}\right)u_{R}(a,\omega)\geq\sum_{\omega\in\Omega}\mu_{0}(\omega)\hat{\sigma}\left(m_{i}|\omega\right)u_{R}(a^{\prime},\omega)\quad\text{for all }a^{\prime}\in A.$
[…]
A.4 Proof of Theorem 1
Here we present and prove a result that generalizes Theorem 1 into a threefold equivalence.
Theorem 1.
Generically, the following statements are equivalent:
  1. (i)
    Commitment is valuable.
  2. (ii)
    Committed Sender values randomization.
  3. (iii)
    For any optimal persuasion profile $(\sigma,\rho)$, there exists $m\in M_{\sigma}$ such that
    $|\arg\max_{a\in A}\sum_{\omega}\mu_{m}(\omega)u_{R}(a,\omega)|\geq 2,$
    where $\mu_{m}$ is defined as $\mu_{m}(\omega)=\frac{\mu_{0}(\omega)\sigma(m|\omega)}{\sum_{\omega}\mu_{0}(\omega)\sigma(m|\omega)}.$
Proof.
We establish the equivalence for any environment that satisfies both partitional-unique-response and scant-indifferences. Since the set of partitional-unique-response environments is generic (Lemma 1) and the set of scant-indifferences environments is generic (Lemma 2), the set of environments that satisfy both properties is also generic.
We will establish that (ii) implies (i), then that (i) implies (iii), and finally that (iii) implies (ii).
Since we are in a scant-indifferences environment, (ii) implies (i) by Proposition 1.
[…]
$\sum_{m}\sigma(m|\omega)u_{S}(\hat{\rho}(m),\omega)\geq u_{S}(\hat{\rho}(m^{\prime}),\omega)$ (12)
[…]
Finally, since we are considering a partitional-unique-response environment, the fact that (iii) implies (ii) is immediate. ∎
A.5 Proof of Theorem 2
Lemma 2 and Proposition 1 jointly imply Theorem 2.
A.6 Proof of Theorem 3
Let $\lambda_{n}$ denote the Lebesgue measure on $\mathbb{R}^{n}$. Recall that the set of environments is $0,1]^{2|A||\Omega|}$. For any property $p$ of an environment, let $\lambda_{2|A||\Omega|}(p):=\lambda_{2|A||\Omega|}(\{(u_{S},u_{R})|(u_{S},u_{R})\text{ satisfies }p\})$ denote the share of environments that satisfy $p$.
Given $u_{S}$, let $\Omega_{i}^{u_{S}}=\{\omega\in\Omega|a_{i}\in\arg\max_{a\in A}u_{S}(a,\omega)\}$ denote the set of states where $a_{i}$ is an ideal action for Sender. [^27] In the body of the paper we denoted this set as $\Omega_{i}$, but for the formal proofs, it is helpful to keep track of the fact that this set depends on $u_{S}$. Note that each $\omega$ must belong to at least one $\Omega_{i}^{u_{S}}$, but the same $\omega$ may appear in multiple $\Omega_{i}^{u_{S}}$. Say that $u_{S}$ is regular if $\Omega_{i}^{u_{S}}\cap\Omega_{j}^{u_{S}}=\emptyset$ for $i\neq j$. We also say that an environment $\left(u_{S},u_{R}\right)$ is regular if $u_{S}$ is regular. Recall that an environment is felicitous if Sender obtains his ideal payoff under cheap talk. We say that an environment is regular-felicitous if it is regular and felicitous. Lemmas 2 and 3 jointly imply that $\lambda_{|A||\Omega|}(\{u_{S}\in[0,1]^{A\times\Omega}|u_{S}\text{ is regular\}})=1$, so $\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{regular-felicity}\right)=\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{felicity}\right)$.
The next lemma provides a convenient characterization of regular-felicity.
Lemma 6.
A regular environment is felicitous if and only if for each $i$,
$\sum_{\omega\in\Omega_{i}^{u_{S}}}\mu_{0}\left(\omega\right)\left(u_{R}\left(a_{i},\omega\right)-u_{R}\left(a^{\prime},\omega\right)\right)\geq 0\quad\text{for all }a^{\prime}\in A.$ (13)
Proof.
Suppose that $u_{S}$ is regular and that (13) holds. Select $|A|$ elements from $M$ and denote them by $m_{1}$ through $m_{|A|}$. Consider the following strategy profile $(\sigma,\rho)$:
  • $\sigma(\omega)=m_{i}$ for every $\omega\in\Omega_{i}^{u_{S}}$;
  • $\rho(m)=a_{i}$ for $m=m_{i}$ ;
  • $\rho(m)=a_{1}$ for $m\notin\{m_{1},...,m_{|A|}\}$.
From regularity and (13), $\left(\sigma,\rho\right)$ is R-BR. In addition, in every state, Sender induces his ideal action, so $\left(\sigma,\rho\right)$ is S-BR and yields Sender’s ideal payoff. Hence, the environment is felicitous.
Conversely, suppose that $u_{S}$ is regular and that Sender obtains his ideal payoff in a cheap-talk equilibrium $(\sigma,\rho)$. By regularity, Sender’s ideal action in each state is unique. This implies that $\rho$ must be pure on-path; otherwise, Sender would fail to induce his unique ideal action with probability 1 in some state.
For each $i$, let $M_{i}=\{m\in M_{\sigma}|\rho(m)=a_{i}\}$ be the set of on-path messages that induce action $a_{i}$. For any $m\in M_{i}$, R-BR implies that
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\sigma(m|\omega)[u_{R}(\omega,a_{i})-u_{R}(\omega,a^{\prime})]\geq 0\quad\text{for all }a^{\prime}\in A.$
Summing the inequality above over $m\in M_{i}$, we obtain
$\sum_{\omega\in\Omega}\sum_{m\in M_{i}}\mu_{0}(\omega)\sigma(m|\omega)[u_{R}(\omega,a_{i})-u_{R}(\omega,a^{\prime})]\geq 0\quad\text{for all }a^{\prime}\in A.$
Since $(\sigma,\rho)$ yields Sender’s ideal payoff, we have that (i) for $\omega\in\Omega_{i}$$\sum_{m\in M_{i}}\sigma(m|\omega)=1$; and (ii) for $\omega\notin\Omega_{i}$, $\sigma(m|\omega)=0$ for all $m\in M_{i}$. Therefore, the inequality simplifies to
$\sum_{\omega\in\Omega_{i}}\mu_{0}(\omega)[u_{R}(\omega,a_{i})-u_{R}(\omega,a^{\prime})]\geq 0\quad\text{for all }a^{\prime}\in A.$
A.6.1 Arbitrary state space
In this section, we establish that for any $\Omega$, the share of environments such that commitment has no value is weakly greater than $\frac{1}{\left|A\right|^{\left|A\right|}}.$
Lemma 7.
$\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{regular-felicity}\right)\geq\frac{1}{\left|A\right|^{\left|A\right|}}$.
Proof.
Fix some regular $u_{S}$. For any non-empty $\Omega_{i}^{u_{S}}$, let
$E_{i}=\{u_{R}\in[0,1]^{A\times\Omega_{i}^{u_{S}}}|\sum_{\omega\in\Omega_{i}^{u_{S}}}\mu_{0}(\omega)u_{R}(a_{i},\omega)>\sum_{\omega\in\Omega_{i}^{u_{S}}}\mu_{0}(\omega)u_{R}(a^{\prime},\omega),\forall a^{\prime}\neq a_{i}\}$
denote the set of Receiver’s preferences on $A\times\Omega_{i}^{u_{S}}$ such that $a_{i}$ is Receiver’s uniquely optimal action given the information that $\omega\in\Omega_{i}^{u_{S}}$. By symmetry, we have that $\lambda_{|A||\Omega_{i}^{u_{S}}|}(E_{i})=\frac{1}{|A|}$. [^28] Note that the set of Receiver’s preferences on $A\times\Omega_{i}^{u_{S}}$ such that Receiver is indifferent between two actions, given the information that $\omega\in\Omega_{i}^{u_{S}}$, has measure zero. The set of $u_{R}\in[0,1]^{A\times\Omega}$ such that $(u_{S},u_{R})$ is felicitous is $\prod_{i:\Omega_{i}^{u_{S}}\text{ is non-empty}}E_{i}$, whose measure is
$\lambda_{\left|A\right|\left|\Omega\right|}\left(\prod_{i:\Omega_{i}^{u_{S}}\text{ is non-empty}}E_{i}\right)$ $=\prod_{i:\Omega_{i}^{u_{S}}\text{ is non-empty}}\lambda_{|A||\Omega_{i}^{u_{S}}|}(E_{i})$
$=\prod_{i:\Omega_{i}^{u_{S}}\text{ is non-empty}}(1/|A|)$
$\geq\prod_{i\in\left\{1,...,\left|A\right|\right\}}(1/|A|)$
$=\frac{1}{|A|^{|A|}}.$ (14)
So, we have established that for any regular $u_{S}$, $\lambda_{\left|A\right|\left|\Omega\right|}(\{u_{R}|(u_{S},u_{R})\text{ is felicitous}\})\geq\frac{1}{|A|^{|A|}}.$ Recall that $\lambda_{\left|A\right|\left|\Omega\right|}(\{u_{S}|u_{S}\text{ is regular}\})=1$. Therefore,
$\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{regular-felicity}\right)$ $=\int_{\{u_{S}|u_{S}\text{ is regular}\}}\int_{\{u_{R}\in[0,1]^{A\times\Omega}|(u_{S},u_{R})\text{ is felicitous}\}}d\lambda_{\left|A\right||\Omega|}\>d\lambda_{\left|A\right|\left|\Omega\right|}$
$\geq\int_{\{u_{S}|u_{S}\text{ is regular}\}}\frac{1}{|A|^{|A|}}d\lambda_{\left|A\right|\left|\Omega\right|}$
$=\frac{1}{|A|^{|A|}}.$
Since commitment has no value in any felicitous environment, Lemma 7 implies that
$\lambda_{2\left|A\right|\left|\Omega\right|}(\text{commitment has no value})$ $\geq\frac{1}{\left|A\right|^{\left|A\right|}}.$
A.6.2 Limit as $\left|\Omega\right|\rightarrow\infty$
In this section, we establish that as $\left|\Omega\right|\rightarrow\infty$: (*) the share of felicitous environments converges to $\frac{1}{\left|A\right|^{\left|A\right|}}$, and (**) generically, commitment has no value if and only if the environment is felicitous. Together, these results imply part (ii) of Theorem 3.
We first give an outline of the proof. We first establish part (**). We show that generically, if the environment is jointly-inclusive, [^29] Recall that an environment is jointly-inclusive if for every action $a$, there is some state $\omega$ such that $a$ is the unique ideal action for both Sender and Receiver in $\omega$. then commitment having no value implies felicity (Lemma 8). We then show that as $\left|\Omega\right|\rightarrow\infty$, the share of jointly-inclusive preferences converges to one (Lemma 9). Combining these two results, we conclude that as $\left|\Omega\right|\rightarrow\infty$, generically, commitment has no value if and only if the environment is felicitous (Lemma 10). To establish part (*), recall that $\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{felicity}\right)\geq\frac{1}{\left|A\right|^{\left|A\right|}}$ and that the reason this is an inequality is the possibility that some $\Omega_{i}^{u_{S}}$ might be empty. We show that as $\left|\Omega\right|\rightarrow\infty$, the share of preferences such that some $\Omega_{i}^{u_{S}}$ is empty converges to zero, which implies that the share of felicitous preferences converges to $\frac{1}{\left|A\right|^{\left|A\right|}}$ (Lemma 11).
Lemma 8.
If commitment has no value in a jointly-inclusive environment that satisfies partitional-unique-response and scant-indifferences, then the environment is felicitous.
Proof.
Consider a jointly-inclusive environment that satisfies partitional-unique-response and scant-indifferences and suppose that commitment has no value. By Proposition 1, there is a partitional $\sigma$ and a pure strategy $\rho$ such that $\left|M_{\sigma}\right|\leq\left|A\right|$ and $(\sigma,\rho)$ is a cheap-talk equilibrium and yields the persuasion payoff.
First, note that every action is induced under $(\sigma,\rho)$; that is, for any $a\in A$, there exists $\omega$ such that $a=\rho(\sigma(\omega))$. To see why, suppose toward contradiction that there is an $a^{*}\in A$ that is not induced. Since the environment is jointly-inclusive, there exists $\omega^{*}$ such that
$u_{S}(a^{*},\omega^{*})>u_{S}(a,\omega^{*})\text{ and }u_{R}(a^{*},\omega^{*})>u_{R}(a,\omega^{*})\text{ for all }a\neq a^{*}.$ (15)
Since $\left|M_{\sigma}\right|\leq\left|A\right|<\left|M\right|$, there is an unsent message, say $m^{*}$.
[…]
  • $\hat{\sigma}(\omega)=\sigma(\omega)$ for $\omega\neq\omega^{*}$, and $\hat{\sigma}\left(m|\omega^{*}\right)=\begin{cases}\left(1-\varepsilon\right)&\text{if }m=\sigma(\omega^{*})\\ \varepsilon&\text{if }m=m^{*}\\ 0&\text{otherwise}\end{cases}$.
  • $\hat{\rho}(m)=\rho(m)$ for $m\neq m^{*}$, and $\hat{\rho}(m^{*})=a^{*}$.
[…]
Now, note that $\rho(\sigma(\omega^{*}))\neq a^{*}$ because $a^{*}$ is not induced under $(\sigma,\rho)$. By (15),
[…]
$>U_{S}(\sigma,\rho).$
This contradicts the fact that $(\sigma,\rho)$ yields the persuasion payoff. Hence, we have established that every action is induced under $(\sigma,\rho)$.
Next, we show that this fact, coupled with the maintained assumptions, implies that the environment is felicitous. Recall that $(\sigma,\rho)$ is a cheap-talk equilibrium; hence for each $\omega$,
$u_{S}(\rho(\sigma(\omega)),\omega)\geq u_{S}(\rho(m),\omega)\text{ for all }m\in M.$
Since every action is induced under $(\sigma,\rho)$, the inequality above is equivalent to
$u_{S}(\rho(\sigma(\omega)),\omega)\geq u_{S}(a,\omega)\text{ for all }a\in A.$
Moreover, since the environment satisfies scant-indifferences, Lemma 3 implies that the environment is regular and
$u_{S}(\rho(\sigma(\omega)),\omega)>u_{S}(a,\omega)\text{ for all }a\neq\rho(\sigma(\omega)).$ (16)
Hence, $\Omega_{i}^{u_{S}}=\{\omega\in\Omega|\rho(\sigma(\omega))=a_{i}\}$ and $\Omega_{i}^{u_{S}}\cap\Omega_{j}^{u_{S}}=\emptyset$ for $i\neq j$. Let $M_{i}=\{m\in M_{\sigma}|\rho(m)=a_{i}\}$. For each $i$ and each $m\in M_{i}$, R-BR of $(\sigma,\rho)$ implies
$\sum_{\omega\in\{\omega:\sigma(\omega)=m\}}\mu_{0}(\omega)u_{R}(a_{i},\omega)\geq\sum_{\omega\in\{\omega:\sigma(\omega)=m\}}\mu_{0}(\omega)u_{R}(a^{\prime},\omega)\text{ for all }a^{\prime}\in A.$
Summing over all $m\in M_{i}$, and noting that $\cup_{m\in M_{i}}\{\omega:\sigma(\omega)=m\}=\{\omega\in\Omega|\rho(\sigma(\omega))=a_{i}\}=\Omega_{i}^{u_{S}}$, we have
$\sum_{\omega\in\Omega_{i}^{u_{S}}}\mu_{0}(\omega)u_{R}(a_{i},\omega)\geq\sum_{\omega\in\Omega_{i}^{u_{S}}}\mu_{0}(\omega)u_{R}(a^{\prime},\omega)\text{ for all }a^{\prime}\in A.$
Hence, by Lemma 6, the environment is felicitous. ∎
Lemma 9.
Fix $A$. As $|\Omega|\rightarrow\infty$, $\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{joint-inclusivity}\right)\rightarrow 1$.
Proof.
For each $\omega\in\Omega$ and $a\in A$, let $E_{a,\omega}^{S}$ denote the set of Sender’s preferences on $A\times\{\omega\}$ such that action $a$ is Sender’s uniquely ideal action in state $\omega$, i.e., $E_{a,\omega}^{S}=\{u_{S}\in\left[0,1\right]^{A\times\left\{\omega\right\}}|u_{S}(a,\omega)>u_{S}(a^{\prime},\omega),\forall a^{\prime}\neq a\}$. Because the set of Sender’s preferences on $A\times\{\omega\}$ such that Sender is indifferent between two actions at $\omega$ has measure zero, by symmetry we have that $\lambda_{\left|A\right|}\left(E_{a,\omega}^{S}\right)=1/|A|$ for each $\left(a,\omega\right)$. Similarly, $\lambda_{\left|A\right|}\left(E_{a,\omega}^{R}\right)=1/|A$ for each $\left(a,\omega\right)$ where $E_{a,\omega}^{R}=\{u_{R}\in[0,1]^{A\times\{\omega\}}|u_{R}(a,\omega)>u_{R}(a^{\prime},\omega),\forall a^{\prime}\neq a\}$. Now, let $E_{a,\omega}=E_{a,\omega}^{S}\times E_{a,\omega}^{R}$ denote the set of preferences on $A\times\{\omega\}$ such that $a$ is the unique ideal action for both Sender and Receiver in state $\omega.$ By the product property of the Lebesgue measure, $\lambda_{2|A|}(E_{a,\omega})=1/|A|^{2}.$
[…]
$\lambda_{2\left|A\right|\left|\Omega\right|}(E_{a})$ $=1-\lambda_{2\left|A\right|\left|\Omega\right|}\left(\prod_{\omega}E_{a,\omega}^{c}\right)$
$=1-\prod_{\omega}\lambda_{2|A|}\left(E_{a,\omega}^{c}\right)$
$=1-\prod_{\omega}\left(1-\frac{1}{|A|^{2}}\right)$
$=1-\left(1-\frac{1}{|A|^{2}}\right)^{|\Omega|}.$
The intersection $\cap_{a\in A}E_{a}$ is the set of jointly-inclusive environments: for every action $a$, there exists at least one state in which $a$ is the unique ideal action for both Sender and Receiver. Therefore,
$\lambda_{2\left|A\right|\left|\Omega\right|}(\text{\text{joint-inclusivity}})=$ $\lambda_{2\left|A\right|\left|\Omega\right|}(\cap_{a\in A}E_{a})$
$=$ $1-\lambda_{2\left|A\right|\left|\Omega\right|}(\cup_{a\in A}E_{a}^{c})$
$\geq$ $1-\sum_{a\in A}\lambda_{2\left|A\right|\left|\Omega\right|}(E_{a}^{c})$
$=$ $1-\sum_{a\in A}\left(1-\lambda_{2\left|A\right|\left|\Omega\right|}(E_{a})\right)$
$=$ $1-\sum_{a\in A}\left(1-\frac{1}{|A|^{2}}\right)^{|\Omega|}$
$=$ $1-|A|\left(1-\frac{1}{|A|^{2}}\right)^{|\Omega|}$
$\rightarrow$ $1\quad\text{as }|\Omega|\rightarrow\infty.$
Lemma 10.
Fix $A$. As $|\Omega|\rightarrow\infty$, generically, commitment has no value if and only if the environment is felicitous.
Proof.
For any $\Omega$, commitment has no value if the environment is felicitous. Hence, we only need to show that as $|\Omega|\rightarrow\infty$, generically, commitment has no value only if the environment is felicitous. Let CNV denote the set of environments in which commitment has no value, and F denote the set of felicitous environments. We need to show that as $|\Omega|\rightarrow\infty$, $\lambda_{2|A||\Omega|}(\text{CNV}\cap\text{F}^{c})\rightarrow 0$.
Let JPS denote the set of environments that are jointly-inclusive and satisfy partitional-unique-response and scant-indifferences. We know from Lemma 8 that $\text{JPS}\cap\text{CNV}\cap\text{F}^{c}=\emptyset$, so $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{JPS}\cap\text{CNV}\cap\text{F}^{c})=0$. As $|\Omega|\rightarrow\infty$, $\lambda_{2\left|A\right|\left|\Omega\right|}\left(\text{JPS}\right)\rightarrow 1$ (Lemmas 1, 2, and 9), which in turn implies $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{JPS}^{c}\cap\text{CNV}\cap\text{F}^{c})\rightarrow 0$. Therefore, as $|\Omega|\rightarrow\infty$, $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{CNV}\cap\text{F}^{c})=\lambda_{2\left|A\right|\left|\Omega\right|}(\text{JPS}\cap\text{CNV}\cap\text{F}^{c})+\lambda_{2\left|A\right|\left|\Omega\right|}(\text{JPS}^{c}\cap\text{CNV}\cap\text{F}^{c})\rightarrow 0.$
Lemma 11.
Fix $A$. As $|\Omega|\rightarrow\infty$, $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{felicity})\rightarrow\frac{1}{|A|^{|A|}}$.
Proof.
Since regular environments are generic, it will suffice to show that $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{regular-felicity})\rightarrow\frac{1}{|A|^{|A|}}$ as $|\Omega|\rightarrow\infty$. Let $E^{S}=\{u_{S}\in[0,1]^{A\times\Omega}|\Omega_{i}^{u_{S}}\text{ is non-empty for all }i\}$ and $E=\{(u_{S},u_{R})\in[0,1]^{(A\times\Omega)^{2}}|u_{S}\in E^{S}\}.$
[…]
$\lambda_{2\left|A\right|\left|\Omega\right|}(\text{regular-felicity}\cap E)$ $=\lambda_{2\left|A\right|\left|\Omega\right|}(\{(u_{S},u_{R})\in[0,1]^{(A\times\Omega)^{2}}|u_{S}\in E^{S},(u_{S},u_{R})\text{ is regular-felicitous})$
$=\int_{E^{S}}\int_{\{u_{R}|(u_{S},u_{R})\text{ is regular-felicitous}\}}d\lambda_{\left|A\right|\left|\Omega\right|}\>d\lambda_{\left|A\right|\left|\Omega\right|}$
$=\int_{E^{S}}\frac{1}{|A|^{|A|}}\,d\lambda_{\left|A\right|\left|\Omega\right|}$
$=\frac{1}{|A|^{|A|}}\lambda_{\left|A\right|\left|\Omega\right|}(E^{S})$
$=\frac{1}{|A|^{|A|}}\lambda_{2\left|A\right|\left|\Omega\right|}(E).$
Note that any jointly-inclusive environment must be contained in $E$. Hence, by Lemma 9, $\lambda_{2\left|A\right|\left|\Omega\right|}(E)\rightarrow 1$ as $|\Omega|\rightarrow\infty.$ Therefore, $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{regular-felicity})\rightarrow\frac{1}{|A|^{|A|}}$ as $|\Omega|\rightarrow\infty$. ∎
Lemmas 10 and 11 jointly yield the fact that, as $|\Omega|\rightarrow\infty$, $\lambda_{2\left|A\right|\left|\Omega\right|}(\text{commitment has no value})\rightarrow\frac{1}{\left|A\right|^{\left|A\right|}}$.
Appendix B Online Appendix
B.1 Role of finite $\Omega$ and AA
In this section, we illustrate some of the issues that arise when $\Omega$ or $A$ is infinite.
B.1.1 Infinite $\Omega$
Suppose $A$ is finite but $\Omega$ is infinite with an atomless prior. It is known (Corollary 1 in zeng2023derandomization) that in this case, randomization is never valuable. Yet, commitment can be valuable. Hence, the analogue of Theorem 1 does not hold.
B.1.2 Infinite AA
Consider a binary-state version of the quadratic-loss, constant-bias preferences from crawford1982strategic. Suppose $\Omega=\left\{0,1\right\}$ and let $\mu_{0}$ be equiprobable. We will contrast the case where $A=\left[0,1\right]$ and the case where $A=\left\{0,\frac{1}{n},\frac{2}{n},...,\frac{n-1}{n},1\right\}$ for some even integer $n\geq 2$. We refer to the former the interval case and the latter as the finite case.
Receiver’s utility is $u_{R}\left(a,\omega\right)=-\left(a-\omega\right)^{2}$. Sender’s utility is $u_{S}\left(a,\omega\right)=-\left(a-\omega-b\right)^{2}$ for some $b>0$. We say a profile $(\sigma,\rho)$ is full revelation if there exist disjoint subsets $M_{0},M_{1}\subseteq M$ such that $\sigma(M_{0}|\omega=0)=1$, $\sigma(M_{1}|\omega=1)=1$.
The parameter region where $b>\frac{1}{2}$ illustrates the contrast between the interval case and the finite case. [^30] When $b\leq\frac{1}{2}$, it does not matter whether we consider the interval or the finite case. In both cases, full revelation yields the persuasion payoff and is a cheap-talk equilibrium. Thus, neither commitment nor randomization is valuable. In the interval case, only a full revelation profile yields the persuasion payoff, but such a profile cannot be a cheap-talk equilibrium. [^31] Due to the strict convexity of Sender’s indirect utility function, the persuasion payoff can only be achieved by full revelation. However, a full revelation profile cannot be a cheap-talk equilibrium: type $\omega=0$ would deviate and send a message in $M_{1},$ because $u_{S}(0,0)=-b^{2}<-(1-b)^{2}=u_{S}(1,0)$. Hence, in the interval case, in contrast to Theorem 1, commitment has value even though committed Sender does not value randomization. In the finite case, however, the assumptions underlying Theorem 1 apply. In this case, commitment is valuable and committed Sender values randomization. [^32] To see that committed Sender values randomization, note that a full revelation profile yields the payoff $\mu_{0}(\omega=1)(0-0-b)^{2}+(1-\mu_{0}(\omega=1))(1-1-b)^{2}]=-b^{2}.$ Providing no information yields the payoff $\mu_{0}(\omega=1)(1/2-0-b)^{2}+(1-\mu_{0}(\omega=1))(1/2-1-b)^{2}]=-(b^{2}+1/4)$. Therefore, Sender’s partitional persuasion payoff is $b^{2}$. But, Sender can obtain a strictly higher payoff by inducing posteriors $\mu=\frac{1}{2n}$ and $\mu=1$. When $\mu=\frac{1}{2n}$, Receiver’s optimal action is $a=\frac{1}{n}$, so Sender’s interim value is $(1-\frac{1}{2n})(\frac{1}{n}-0-b)^{2}+(\frac{1}{2n})(\frac{1}{n}-1-b)^{2}]=-b^{2}+\frac{2b-1}{2n}>-b^{2}$. When $\mu=1$, Sender’s interim value is $b^{2}$. Sender’s ex-ante payoff is a convex combination of the two interim values, which is strictly higher than $b^{2}.$
B.2 State-independent preferences
As we mentioned in the discussion of related literature, several papers examine value of commitment under the assumption that Sender has state-independent preferences. To connect to that literature, it is worth asking whether our results hold under that assumption. When $\left|A\right|\geq 3$, state-independent preferences by Sender mean that we are not in a scant-indifferences environment, so the proofs above do not apply. Nonetheless, Theorems 1 and 2 indeed remain true.
To formalize this, say that environment $\left(u_{S},u_{R}\right)$ is transparent if there exists some function $v:A\rightarrow\mathbb{R}$ such that $u_{S}\left(a,\omega\right)=u_{S}(a,\omega^{\prime})\equiv v(a)$ for any $a,\omega,\omega^{\prime}$. The set of all transparent environments is $0,1]^{\left|A\right|\,\left(\left|\Omega\right|+1\right)}$. A set of transparent environments is transparently-generic if it has Lebesgue measure one in $\mathbb{R}^{\left|A\right|\,\left(\left|\Omega\right|+1\right)}$. We say that a claim holds generically in transparent environments, if it holds for a transparently-generic set of environments.
A transparent environment $\left(u_{S},u_{R}\right)$ satisfies no-duplicate-actions if for any $a\neq a^{\prime}$, $v(a)\neq v(a^{\prime})$. Clearly, the set of no-duplicate-actions transparent environments is transparently-generic.
A strategy profile $(\sigma,\rho)$ is a simple babbling cheap-talk equilibrium if it is a cheap-talk equilibrium in which $|M_{\sigma}|=1$ and $\rho(m)=a_{0}$ for all $m\in M$ and for some $a_{0}\in A$.
Lemma 12.
In a no-duplicate-actions transparent environment, if commitment has no value, then there exists a simple babbling cheap-talk equilibrium that yields the persuasion payoff.
Proof.
If commitment has no value, some cheap-talk equilibrium, denoted by $\left(\sigma,\rho\right)$, yields the persuasion payoff. First, we claim that $\rho$ must be pure on-path. Suppose by contradiction that at some on-path message $m$, $|\operatorname{Supp}(\rho(\cdot|m))|>1$. R-BR then implies that Receiver must be indifferent among all actions in $\operatorname{Supp}(\rho(\cdot|m))$. Since the environment satisfies no-duplicate-actions, Sender must strictly prefers one of the actions in $\operatorname{Supp}(\rho(\cdot|m))$ over all others. Therefore, an alternative strategy profile where Receiver breaks ties in favor of Sender would still satisfy R-BR while strictly improving Sender’s payoff.
Since $\left(\sigma,\rho\right)$ is a cheap-talk equilibrium and the environment is transparent, S-BR implies that for any $m,m^{\prime}\in M_{\sigma}$, we have $v(\rho(m))=v(\rho(m^{\prime}))$. Moreover, because the environment satisfies no-duplicate-actions, it follows that $\rho(m)=\rho(m^{\prime})$ for any $m,m^{\prime}\in M_{\sigma}$; that is, only a single action is induced in equilibrium.
[…]
$\sum_{\omega\in\Omega}\mu_{0}(\omega)\sum_{m\in M}\sigma(m|\omega)u_{R}(\rho(m),\omega)\geq\sum_{\omega\in\Omega}\mu_{0}(\omega)\sum_{m\in M}\sigma(m|\omega)u_{R}(a,\omega)\quad\text{for all }a\in A.$
Since $\rho(m)=\rho(m_{0})$ for all $m\in M_{\sigma}$, we can rewrite the inequality as
$\sum_{\omega\in\Omega}\mu_{0}(\omega)u_{R}(\rho(m_{0}),\omega)\geq\sum_{\omega\in\Omega}\mu_{0}(\omega)u_{R}(a,\omega)\quad\text{for all }a\in A,$
[…]
Theorem 4.
Generically in transparent environments, commitment is valuable if and only if committed Sender values randomization.
Proof.
We establish the equivalence for any transparent environment that satisfies both partitional-unique-response and no-duplicate-actions. Recall that whether an environment satisfies partitional-unique-response does not depend on Sender’s preferences, so the same argument as in Lemma 1 implies that the set of partitional-unique-response transparent environments is transparently-generic. Moreover, the set of no-duplicate-actions transparent environments is transparently-generic. Therefore, the set of transparent environments that satisfy both properties is also transparently-generic.
The same arguments that establish (i) implies (iii) and (iii) implies (ii) in Theorem 1’ apply directly to any partitional-unique-response transparent environment, thereby proving the “only if” direction. The “if” direction follows immediately from Lemma 12. ∎
Theorem 5.
Generically in transparent environments, commitment is valuable if cheap-talk Sender values randomization.
Proof.
The theorem follows immediately from Lemma 12 and the fact that the set of no-duplicate-actions transparent environments is transparently-generic. ∎
B.3 Value of partial commitment
We consider a partial commitment setting where Sender can commit to a distribution of messages, as in lin2024credible. For any Sender’s strategy $\sigma,$ let $D(\sigma):=\{\sigma^{\prime}:\Omega\rightarrow\Delta M|\sum_{\omega}\mu_{0}(\omega)\sigma^{\prime}(m|\omega)=\sum_{\omega}\mu_{0}(\omega)\sigma(m|\omega),\forall m\}$ denote the set of messaging strategies that preserve the same distribution of messages.
We say a profile $(\sigma^{*},\rho^{*})$ is a curve equilibrium if
$\sigma^{*}\in\operatorname*{arg\,max}_{\sigma\in D(\sigma^{*})}U_{S}(\sigma,\rho^{*})$
$\rho^{*}\in\operatorname*{arg\,max}_{\rho}U_{R}(\sigma^{*},\rho).$
The first condition requires that Sender has no incentive to deviate to any other messaging strategy that preserves the same distribution of messages as $\sigma^{*}$, and the second condition requires Receiver to play a best response (i.e., R-BR).
The curve payoff is the maximum $U_{S}$ induced by a curve equilibrium. The curve partitional payoff is the maximum $U_{S}$ induced by a partitional curve equilibrium. We say Sender values committing to a curve if the curve payoff is strictly higher than the cheap-talk payoff, and that curve-committed Sender values randomization if the curve payoff is strictly higher than the curve partitional payoff.
A function $u_{S}:A\times\Omega\rightarrow\mathbb{R}$ is strictly supermodular if there exists a total order $_{A}$ on $A$ and a total order $_{\Omega}$ on $\Omega$ such that for $a^{\prime}>_{A}a$ and $\omega^{\prime}>_{\Omega}\omega$,
$u_{S}(a^{\prime},\omega^{\prime})-u_{S}(a,\omega^{\prime})>u_{S}(a^{\prime},\omega)-u_{S}(a,\omega).$
To simplify notation, once we fix a strictly supermodular $u_{S}$, we use >> in place of $_{A}$ and $_{\Omega}$.
A set of Receiver’s preferences is generic if it has Lebesgue measure one in $0,1]^{\left|A\right|\,\left|\Omega\right|}$.
We say that an outcome distribution $\pi:\Omega\rightarrow\Delta A$ is induced by a profile $(\sigma,\rho)$ if $\pi(a|\omega)=\sum_{m\in M}\rho(a|m)\sigma(m|\omega)$ for all $\omega,a$. The following lemma offers a characterization of outcome distributions that can be induced by a curve equilibrium.
Lemma 13.
Fix any strictly supermodular $u_{S}$. An outcome distribution $\pi:\Omega\rightarrow\Delta A$ is induced by some curve equilibrium $(\sigma,\rho)$ where $\rho$ is pure strategy on-path if and only if
[…]
2. $\pi$ is $u_{R}$-obedient: for each $a,a^{\prime}\in A$,
$\sum_{\omega\in\Omega}\pi(a|\omega)\mu_{0}(\omega)\left[u_{R}(a,\omega)-u_{R}(a^{\prime},\omega)\right]\geq 0.$
Proof.
The lemma follows immediately from Theorem 1 and Lemma 1 in lin2024credible, with the small caveat that in lin2024credible, $\rho$ is restricted to be pure both on and off-path. However, note that off-path actions do not affect whether a profile is a curve equilibrium. Thus, the lemma follows. ∎
Theorem 6.
Fix any strictly supermodular $u_{S}$. For a generic set of Receiver’s preferences, if Sender values committing to a curve, then a curve-committed Sender values randomization.
Proof.
We consider any Receiver preference that satisfies partitional-unique-response. By Lemma 1, this set of preferences is generic.
[…]
$u_{S}(\hat{\rho}(\sigma(\omega)),\omega)\geq u_{S}(\hat{\rho}(m^{\prime}),\omega)$ (17)
[…]
Since $(\sigma,\rho)$ is a curve equilibrium and $\rho$ is pure on-path, by Lemma 13, the induced outcome distribution, denoted by $\pi$, satisfies comonotonicity and $u_{R}$-obedience. In addition, since $\sigma$ is partitional and $\rho$ is pure on-path, the induced outcome distribution $\pi$ is also a pure mapping. Moreover, $\pi$ being comonotone can be strengthened to $\pi$ being monotone partitional:
$\forall a<a^{\prime},\text{ }\pi(a|\omega)>0\text{ and }\pi(a^{\prime}|\omega^{\prime})>0\text{ implies }\omega<\omega^{\prime}.$ (18)
Moreover, since the environment satisfies partitional-unique-response, $u_{R}$-obedience can be strengthened to strict $u_{R}$-obedience: for each $a\in A^{*}\equiv\cup_{\omega\in\Omega}\operatorname{Supp}(\pi(\cdot|\omega))$ and $a^{\prime}\in A/\{a\}$,
$\sum_{\omega\in\Omega}\pi(a|\omega)\mu_{0}(\omega)\left[u_{R}(a,\omega)-u_{R}(a^{\prime},\omega)\right]>0.$ (19)
For each $a\in A^{*}$, let $\Omega_{a}=\{\omega|\hat{\rho}(\sigma(\omega))=a\}$ denote the set of states that induce action $a$. By (18), $\{\Omega_{a}\}_{a\in A^{*}}$ forms a monotone partition of $\Omega$: $\cup_{a\in A^{*}}\Omega_{a}=\Omega$, and for any $a<a^{\prime}$, $\omega\in\Omega_{a}$, $\omega^{\prime}\in\Omega_{a^{\prime}}$, we have $\omega<\omega^{\prime}$.
[…]
Let $a^{*}\equiv\pi(\omega^{*})$ and $\hat{a}\in\min\{\operatorname*{arg\,max}_{a^{\prime}>a^{*},a^{\prime}\in A^{*}}u_{S}(a^{\prime},\omega)\}$ denote type $\omega^{*}$’s smallest optimal action among $\{a^{\prime}|a^{\prime}>a^{*}\}$. Let $\tilde{\omega}=\max\{\omega|\pi(\omega)<\hat{a}\}$ denote the largest type that induces an action smaller than $\hat{a}$. Let $\tilde{a}\equiv\pi(\tilde{\omega})<\hat{a}$. Since $\hat{a}$ is $\omega^{*}$’s smallest optimal action among $\{a^{\prime}|a^{\prime}>a^{*}\}$, $u_{S}(\hat{a},\omega^{*})>u_{S}(\tilde{a},\omega^{*}).$ By definition, $\tilde{\omega}\geq\omega^{*}$, so by supermodularity,
$u_{S}(\hat{a},\tilde{\omega})-u_{S}(\tilde{a},\tilde{\omega})\geq u_{S}(\hat{a},\omega^{*})-u_{S}(\tilde{a},\omega^{*})>0.$ (20)
[…]
Since $\hat{\pi}$ that satisfies comonotonicity and $u_{R}$-obedience, by Lemma 13, there exists a curve equilibrium that yields a strictly higher payoff than $(\sigma,\rho)$. This contradicts the fact that $(\sigma,\rho)$ yields the curve payoff. ∎
B.4 Away from zeros
Theorem 1 tells us that, generically, commitment has zero value if and only if randomization has zero value. A natural question is whether, generically, a small value of commitment implies or is implied by a small value of randomization. This section establishes that the answer is no.
Specifically, we show that there is a $\Delta>0$ such that for any $\varepsilon>0$: (i) there is a positive measure of environments where the value of commitment is greater than $\Delta$ and the value of randomization is less than $\varepsilon$, and (ii) there is a positive measure of environments where the value of randomization is greater than $\Delta$ and the value of commitment is less than $\varepsilon$. [^33] Recall that we assume that the range of the utility functions is $\left[0,1\right]$ in order to succinctly state our results about how often commitment is valuable. If we relax the assumption that the range of the utility functions is uniformly bounded, we can alternatively show that for any $\Delta>0$ and any $\varepsilon>0$: (i) there is a positive measure of environments where the value of commitment is greater than $\Delta$ and the value of randomization is less than $\varepsilon$, and (ii) there is a positive measure of environments where the value of randomization is greater than $\Delta$ and the value of commitment is less than $\varepsilon$.
We begin by illustrating the role of the genericity condition for Theorem 1. We present two examples. The first example presents an environment (that violates partitional-unique-response) where commitment is valuable but randomization is not. The second example presents an environment (that violates scant indifferences) where randomization is valuable but commitment is not.
Then, we build on the first example to show to construct, given any $\varepsilon$, a positive measure of environments where the value of commitment is above $\frac{1}{6}$ but the value of randomization is less than $\varepsilon$. We build on the second example to show how to construct, given any $\varepsilon$, a positive measure of environments where the value of randomization is above $\frac{1}{13}$ but the value of of commitment is less than $\varepsilon$.
Example 1.
Consider $\Omega=\{\omega_{1},\omega_{2}\}$ with prior $\mu_{0}(\omega_{1})=\mu_{0}(\omega_{2})=0.5$ and $A=\{a_{1},a_{2}\}$. Players’ payoffs are given in Table 1. Receiver’s best response is $a_{1}$ when $\mu\equiv\mu(\omega_{2})\in[0,1)$, and she is indifferent between $a_{1}$ and $a_{2}$ when $\mu=1$. The concavification of Sender’s indirect utility function is depicted in Figure 2.
$u_{S}$ $a_{1}$ $a_{2}$
$\omega_{1}$ 0 $2/3$
$\omega_{2}$ 0 $2/3$
$u_{R}$ $a_{1}$ $a_{2}$
$\omega_{1}$ 1/2 0
$\omega_{2}$ 1/2 1/2
[…]
Table 1: Sender and Receiver’s payoffs
Figure 2: Concavification
Clearly, full revelation is the unique optimal information structure, which yields a payoff of $1/3$; therefore, Sender does not value randomization. In addition, the only possible cheap-talk equilibrium outcome is babbling, which yields a payoff of 0. Hence, commitment is valuable.
Example 2.
Consider $\Omega=\{\omega_{1},\omega_{2}\}$ with prior with prior $\mu_{0}(\omega_{1})=\mu_{0}(\omega_{2})=0.5$ and $A=\{a_{1},a_{2},a_{3}\}$. Players’ payoffs are given in Table 2. Receiver’s best response is $a_{1}$ when $\mu\in[0,1/3]$, $a_{2}$ when $\mu\in[1/3,2/3]$, and $a_{3}$ when $\mu\equiv\mu(\omega_{2})\in[2/3,1]$. This leads to Sender’s indirect utility function (blue) and its concave envelope (red) depicted in Figure 3.
$u_{S}$ $a_{1}$ $a_{2}$ $a_{3}$
$\omega_{1}$ $1/3$ $2/3$ 0
$\omega_{2}$ $5/6$ $1/6$ $1/2$
$u_{R}$ $a_{1}$ $a_{2}$ $a_{3}$
$\omega_{1}$ $3/4$ $1/2$ 0
$\omega_{2}$ 0 $1/2$ $3/4$
[…]
Table 2: Sender’s and Receiver’s payoffs
Figure 3: Concavification
Sender values randomization, because the unique optimal information structure that induces a posterior $1/3$ cannot be generated by a partitional messaging strategy. Since Sender’s indirect utility function is continuous, by lipnowski2020equivalence, Sender does not value commitment. For the sake of completeness, we construct a cheap-talk equilibrium that yields the persuasion payoff $1/2$ to Sender.
Consider a strategy profile $(\sigma,\rho)$ with two on-path messages $m_{1},m_{2}$: $\sigma(m_{1}|\omega_{2})=1/2$, $\sigma(m_{2}|\omega_{2})=1/2$, $\sigma(m_{1}|\omega_{1})=1$; $\rho(a_{1}|m_{1})=1/2$, $\rho(a_{2}|m_{1})=1/2$ $\rho(a_{3}|m_{2})=1.$
The profile satisfies R-BR because the posterior upon observing $m_{1}$ is 1/3 and when observing $m_{2}$ is $1$. We next show that the profile also satisfies S-BR. For type $\omega_{2}$ Sender, the expected payoff of sending message $m_{2}$ is 1/2 and the expected payoff of sending message $m_{1}$ is $\frac{1}{2}\cdot\frac{5}{6}+\frac{1}{2}\cdot\frac{1}{6}=\frac{1}{2}$, so type $\omega_{2}$ Sender is indifferent and has no incentive to deviate. For type $\omega_{1}$ Sender, the expected payoff of sending message $m_{2}$ is 0 and the expected payoff of sending message $m_{1}$ is $\frac{1}{2}\cdot\frac{1}{3}+\frac{1}{2}\cdot\frac{2}{3}=\frac{1}{2}$, so he strictly prefer sending message $m_{1}$. Hence, $(\sigma,\rho)$ is a cheap-talk equilibrium that yields the persuasion payoff.
B.4.1 Small value of randomization does not imply small value of commitment
We construct a positive measure of environments where the value of commitment is bounded away from zero but the value of randomization is arbitrarily small. Formally, given an environment $(u_{S},u_{R})$, let $\Delta_{\mathcal{R}}(u_{S},u_{R})=\text{\emph{Persuasion Payoff}}-\text{\emph{Partitional Persuasion Payoff}}$ and $\Delta_{\mathcal{C}}(u_{S},u_{R})=\text{\emph{Persuasion Payoff}}-\text{\emph{Cheap-Talk Payoff}}$ denote the value of randomization and the value of commitment, respectively. We will establish that for any $\varepsilon>0$, there exists a positive measure set of environments $E$ such that for any $(u_{S},u_{R})\in E$, we have $\Delta_{\mathcal{R}}(u_{S},u_{R})<\varepsilon$ and $\Delta_{\mathcal{C}}(u_{S},u_{R})>1/6$.
Fix any $\varepsilon>0$, we perturb players’ payoffs in Example 1 to construct the desired positive share of environments. The idea is that for sufficiently small perturbations in an appropriate direction, the changes to the persuasion payoff, partitional persuasion payoff, and cheap-talk payoff will also be small. Since in Example 1, the value of randomization is zero and the value of commitment is positive, we obtain a positive measure of environments with a small value of randomization and a value of commitment bounded away from zero.
Players’ payoffs are as in Table 3, where $s_{ij},r_{ij}$ are the perturbations to Sender’s and Receiver’s payoffs, respectively, when action $a_{j}$ is taken in state $\omega_{i}$.
$u_{S}$ $a_{1}$ $a_{2}$
$\omega_{1}$ $0+s_{11}$ $2/3+s_{12}$
$\omega_{2}$ $0+s_{21}$ $2/3+s_{22}$
$u_{R}$ $a_{1}$ $a_{2}$
$\omega_{1}$ $_{11} r12r_{12}$
$\omega_{2}$ $1/2+r_{21}$ $1/2+r_{22}$
Table 3: Sender’s and Receiver’s payoffs
Let $s_{ij}\in[0,\delta]$, $r_{11},r_{21}\in[-\delta,0]$, and $r_{12},r_{22}\in[0,\delta]$, where $\delta>0$. These perturbations generate a positive measure set of environments, denoted by $E^{\delta}$. We will establish that if $\delta<\min\{\frac{1}{4},\frac{\varepsilon}{2(1+\varepsilon)}\},$ then for any $(u_{S},u_{R})\in E^{\delta},$ the value of commitment is greater than $1/6$ and the value of randomization is less than $\varepsilon$.
Consider any $(u_{S},u_{R})\in E^{\delta}$. Since $\delta<1/4,$ Receiver’s best response is $a_{2}$ iff $\mu\geq\mu^{*}\equiv\frac{1/2+r_{11}-r_{12}}{1/2+r_{11}-r_{21}-r_{12}+r_{22}}\in[1-2\delta,1].$ Sender’s indirect utility function is therefore
$v(\mu)=\begin{cases}s_{11}+(s_{21}-s_{11})\mu&\text{if }\mu<\mu^{*}\\ \frac{2}{3}+s_{12}+(s_{22}-s_{12})\mu&\text{if }\mu\geq\mu^{*}.\end{cases}$
By inducing beliefs $\mu=0$ and $\mu=\mu^{*}$, Sender achieves her persuasion payoff
$\frac{1}{2\mu^{*}}[\frac{2}{3}+s_{12}+(s_{22}-s_{12})\mu^{*}]+(1-\frac{1}{2\mu^{*}})s_{11}.$
Meanwhile, full revelation yields a payoff of
$\frac{1}{3}+\frac{s_{11}+s_{22}}{2}.$
Therefore,
$\Delta_{\mathcal{R}}(u_{S,}u_{R})\leq$ $\frac{1}{2\mu^{*}}[\frac{2}{3}+s_{12}+(s_{22}-s_{12})\mu^{*}]+(1-\frac{1}{2\mu^{*}})s_{11}-\frac{1}{3}-\frac{s_{11}+s_{22}}{2}$
$=$ $\frac{1-\mu^{*}}{\mu^{*}}(\frac{1}{3}+\frac{s_{12}-s_{11}}{2})$
$\leq$ $\frac{2\delta}{1-2\delta}(\frac{1}{3}+\delta)$
$<$ $\varepsilon$
where the third line follows from $\mu^{*}\geq 1-2\delta$ and $s_{ij}\in[0,\delta]$, and the last line follows from $\delta<\frac{\varepsilon}{2(1+\varepsilon)}$ and $\delta<\frac{2}{3}$.
In addition, since $\delta<1/4$, $a_{2}$ is Sender’s preferred action regardless of the states. It follows that any cheap-talk equilibrium outcome must be the babbling outcome where Receiver takes action $a_{1}$ with probability 1. Therefore, Sender’s cheap-talk payoff is $\frac{s_{11}+s_{21}}{2}$. Hence, $\Delta_{\mathcal{C}}(u_{S,}u_{R})=\text{\emph{Persuasion Payoff}}-\text{\emph{Cheap-Talk Payoff}}\geq\frac{1}{3}+\frac{s_{11}+s_{22}}{2}-\frac{s_{11}+s_{21}}{2}>\frac{1}{6}$, where the weak inequality follows from the fact that the persuasion payoff is greater than the payoff from full revelation, and the strict inequality follows from $\frac{s_{22}-s_{21}}{2}\geq\frac{-\delta}{2}>-\frac{1}{6}$.
B.4.2 Small value of commitment does not imply small value of randomization
We will establish that for any $\varepsilon>0$, there exists a positive measure set of environments $E$ such that for any $(u_{S},u_{R})\in E$, we have $\Delta_{\mathcal{C}}(u_{S},u_{R})<\varepsilon$ and $\Delta_{\mathcal{R}}(u_{S},u_{R})>\frac{1}{13}$.
Fix any $\varepsilon>0$, we perturb players’ payoffs in Example 2 to construct the desired positive share of environments. Similar to Section B.4.1, the idea is to show that changes to the persuasion payoff, partitional persuasion payoff, and cheap-talk payoff are small under small perturbations. Since in Example 2, the value of commitment is zero and the value of randomization is positive, we obtain a positive measure of environments with a small value of commitment and a value of randomization bounded away from zero.
Players’ payoffs are as in Table 4, where $s_{ij},r_{ij}$ are the perturbations to Sender’s and Receiver’s payoffs, respectively, when action $a_{j}$ is taken in state $\omega_{i}$.
$u_{S}$ $a_{1}$ $a_{2}$ $a_{3}$
$\omega_{1}$ $1/3+s_{11}$ $2/3+s_{12}$ $0+s_{13}$
$\omega_{2}$ $5/6+s_{21}$ $1/6+s_{22}$ $1/2+s_{23}$
$u_{R}$ $a_{1}$ $a_{2}$ $a_{3}$
$\omega_{1}$ $3/4+r_{11}$ $1/2+r_{12}$ $0+r_{13}$
$\omega_{2}$ $0+r_{21}$ $1/2+r_{22}$ $3/4+r_{23}$
Table 4: Sender’s and Receiver’s payoffs
Let $s_{ij},r_{ij}\in[0,\delta]$, where $\delta>0$. These perturbations generate a positive measure set of environments, denoted by $E^{\delta}$. Consider any $(u_{S},u_{R})\in E^{\delta}$ with $\delta<1/4$. Receiver’s best response is
$a_{R}(\mu)=\begin{cases}a_{1}&\text{if }\mu\in[0,\mu_{12}]\\ a_{2}&\text{if }\mu\in[\mu_{12},\mu_{23}]\\ a_{3}&\text{if }\mu\in[\mu_{23},1]\end{cases}$
where $\mu_{12}=\frac{1}{3+r_{12}+r_{22}-r_{12}-r_{21}}$ and $\mu_{23}=\frac{2+r_{12}-r_{13}}{3+r_{23}-r_{13}+r_{12}-r_{22}}.$ Similar to Example 2 , the optimal information structure induces beliefs $\mu_{12}$ and $1$, yielding a value
$\frac{1}{2(1-\mu_{12})}\max\{\mu_{12}(5/6+s_{21})+(1-\mu_{12})(1/3+s_{11}),\mu_{12}(1/6+s_{22})+(1-\mu_{12})(2/3+s_{12})\}\\ +\frac{1-2\mu_{12}}{2(1-\mu_{12})}(1/2+s_{23}).$
Since $s_{ij}\in[0,\delta],$ taking $\delta\rightarrow 0$, we have $\mu_{12}\rightarrow 1/3$, and the above value approaches $1/2$. By continuity, there exists $\delta^{1}>0$ such that for any $\delta<\delta^{1}$, the persuasion value lies within the interval $\frac{1}{2}-\frac{\varepsilon}{2},\frac{1}{2}+\frac{\varepsilon}{2}].$
Meanwhile, full revelation yields a payoff of $\frac{5/6+s_{11}+s_{23}}{2}$ and providing no information yields a payoff of $\frac{5/6+s_{12}+s_{22}}{2}$. Both values approach $5/12$ when $\delta\rightarrow 0$. By continuity, there exists $\delta^{2}>0$ such that for any $\delta<\delta^{2}$, the partitional persuasion value is less than $\frac{5}{12}+\frac{\varepsilon}{2}$.
Next, we will construct a cheap-talk equilibrium that yields a payoff close to $1/2$ for small $\delta$. Consider a strategy profile $(\sigma,\rho)$ with two on-path messages $m_{1},m_{2}$: $\sigma(m_{1}|\omega_{2})=\frac{\mu_{12}}{1-\mu_{12}}$, $\sigma(m_{2}|\omega_{2})=\frac{1-2\mu_{12}}{1-\mu_{12}}$, $\sigma(m_{1}|\omega_{1})=1$; $\rho(a_{1}|m_{1})=p$, $\rho(a_{2}|m_{1})=1-p$, $\rho(a_{3}|m_{2})=1$, where $p=\frac{1/3+s_{23}-s_{22}}{2/3+s_{21}-s_{22}}$. Since $\delta<1/3$, $p\in(0,1)$ is a well-defined probability.
The strategy profile satisfies R-BR because the posterior upon observing $m_{1}$ is $\mu_{12}$, and upon observing $m_{2}$ is $1$. We now show that the profile also satisfies S-BR. For type $\omega_{2}$ Sender, the expected payoff of sending message $m_{2}$ is $1/2+s_{23}$ and the expected payoff of sending message $m_{1}$ is $p(5/6+s_{21})+(1-p)(1/6+s_{22})=1/2+s_{23}$, so type $\omega_{2}$ Sender is indifferent and has no incentive to deviate. For type $\omega_{1}$ Sender, the expected payoff of sending message $m_{2}$ is $0+s_{13}$ and the expected payoff of sending message $m_{1}$ is $p(1/3+s_{11})+(1-p)(2/3+s_{12})$. As $\delta\rightarrow 0$, $p(1/3+s_{11})+(1-p)(2/3+s_{12})\rightarrow 1/2$ and $0+s_{13}\rightarrow 0$, so type $\omega_{1}$ Sender strictly prefers to send message $m_{1}$. By continuity, there exists $\delta^{3}>0$ such that for any $\delta<\delta^{3},$ the strategy profile $(\sigma,\rho)$ is a cheap-talk equilibrium, yielding a value $\frac{1}{2}(1/2+s_{23})+\frac{1}{2}[p(1/3+s_{11})+(1-p)(2/3+s_{12})]>1/2-\frac{\varepsilon}{2}$.
Therefore, for any $\delta<\delta^{*}\equiv\min\{\delta^{1},\delta^{2},\delta^{3},\frac{1}{4}\}$, and for any $(u_{S},u_{R})\in E^{\delta}$, $\Delta_{\mathcal{C}}\left(u_{S},u_{R}\right)<(\frac{1}{2}+\frac{\varepsilon}{2})-(\frac{1}{2}-\frac{\varepsilon}{2})=\varepsilon$, and $\Delta_{\mathcal{R}}\left(u_{S},u_{R}\right)>(\frac{1}{2}-\frac{\varepsilon}{2})-(\frac{5}{12}+\frac{\varepsilon}{2})=\frac{1}{12}-\varepsilon$.
We have established that for any $\varepsilon>0$, there exists a positive measure set of environments $E$ such that for any $(u_{S},u_{R})\in E$, we have $\Delta_{\mathcal{C}}(u_{S},u_{R})<\varepsilon$ and $\Delta_{\mathcal{R}}(u_{S},u_{R})>\frac{1}{12}-\varepsilon$. Now fix any $\hat{\varepsilon}>0$, let $\varepsilon=\min\{\hat{\varepsilon},\frac{1}{12}-\frac{1}{13}\}$, our previous result implies that there exists a positive measure set of environments $E$ such that for any $(u_{S},u_{R})\in E$, $\Delta_{\mathcal{C}}(u_{S},u_{R})<\varepsilon\leq\hat{\varepsilon}$ and $\Delta_{\mathcal{R}}(u_{S},u_{R})>\frac{1}{12}-\varepsilon\geq\frac{1}{12}-(\frac{1}{12}-\frac{1}{13})=\frac{1}{13}$.